### v1

# Look before Transcription: End-to-End SlideASR with Visually-Anchored Policy Optimization

Abstract
Automatic speech recognition (ASR) systems often struggle with domain-specific terminology, especially in specialized settings such as academic lectures. To address this, we define the SlideASR task, which leverages the rich visual information from presentation slides to improve transcription accuracy. Existing pipeline methods for this task tend to be complex and underperform. Although omni-modal large language models (OLLMs) provide a promising end-to-end framework, they frequently fail in practice by degenerating into simple optical character recognition (OCR) systems. To overcome this, we propose Visually-Anchored Policy Optimization (VAPO), a novel post-training method designed to control the model’s reasoning process. Drawing on the Chain-of-Thought reasoning paradigm, VAPO enforces a structured "Look before Transcription" procedure using a <think><answer> format. Specifically, the model first performs OCR on the slide content within the think step, then generates the transcription by referencing this recognized visual information in the answer step. This reasoning process is optimized via reinforcement learning with four distinct rewards targeting format compliance, OCR accuracy, ASR quality, and visual anchoring consistency. To support further research, we construct SlideASR-Bench, a new entity-rich benchmark consisting of a synthetic dataset for training and testing, and a challenging real-world set for evaluation. Extensive experiments demonstrate that VAPO significantly improves recognition of domain-specific terms, establishing an effective end-to-end paradigm for SlideASR.
Look before Transcription: End-to-End SlideASR with
Visually-Anchored Policy Optimization
[…]
1 Introduction
Current end-to-end automatic speech recognition (ASR) models, such as Whisper Radford et al. (2023), have demonstrated impressive performance in transcribing common words. However, recognition performance often deteriorates significantly in specialized domains, such as academic lectures, technical presentations, or medical seminars. Previous works have improved ASR accuracy by incorporating lip movement information from the speaker Afouras et al. (2022); Ma et al. (2021, 2023); Shi et al. (2022). However, in addition to facial information, there is multimodal textual information on the slide that is closely related to the current speech of the speaker Wang et al. (2024b); Zhao et al. (2025), which lipreading-based approaches cannot utilize. For clarity, we refer to the task of improving ASR accuracy by incorporating visual cues from presentation slides as SlideASR.
Figure 1: We compare the outputs of Qwen2.5-Omni-7B, Qwen3-Omni-30B-A3B (with OCR text as context), and our VAPO-7B on a real Chinese medical report sample. Red text indicates incorrectly transcribed named entities.
Currently, the dominant strategy for SlideASR task is the pipeline paradigm. Specifically, these methods can be divided into two categories. The first is post-processing correction Trinh et al. (2025), which utilizes large language models (LLMs) for text-level refinement after independent ASR and optical character recognition (OCR). The second is contextual enhancement Wang et al. (2024b, a); Yu et al. (2024); Yang et al. (2024), where OCR-extracted text is injected as context into the ASR model. The post-processing correction methods involve multiple modules, resulting in a complex workflow and relatively high latency. In contrast, for contextual enhancement methods, we observed that when OCR text is provided to large audio language models (LALMs) Chu et al. (2024); Dinkel et al. (2025), the models often tend to repeat the context rather than effectively assisting the recognition process.
Beyond the pipeline paradigm, we argue that the recently emerging omni-modal large language models (OLLMs) Xu et al. (2025a, b); Yao et al. (2024); Li et al. (2025a) are inherently well-suited for the SlideASR task. OLLMs are capable of simultaneously processing textual, visual, and auditory modalities, and performing tasks based on human instructions. Therefore, they are theoretically able to accomplish the SlideASR task in an end-to-end manner. However, we identify two key issues with existing OLLMs when applied to the SlideASR task. First, their behavior is often uncontrollable. When given an audio clip and the corresponding slide image, the models do not consistently perform accurate transcription. In some cases, they simply copy text from the slide, effectively acting as OCR systems instead of transcribing speech. Second, even when ASR is successful, the model often fails to leverage entity information from the slide. Entities that are clearly shown in the image are sometimes transcribed incorrectly, suggesting weak integration of visual context. These issues highlight that the implicit and opaque reasoning processes of OLLMs are unreliable for complex multimodal tasks such as SlideASR. Inspired by recent advances in LLMs, where explicit Chain-of-Thought (CoT) reasoning Guo et al. (2025); Jaech et al. (2024); Shao et al. (2024b) has been shown to improve performance on complex problems, we argue that incorporating structured reasoning into the SlideASR task can help constrain model behavior and address the challenges of multimodal fusion.
To this end, we propose Visually-Anchored Policy Optimization (VAPO), a post-training method tailored to enhance the performance of OLLMs on the SlideASR task. The core idea of VAPO is to replace the model’s implicit and uncontrolled reasoning with an explicit, structured process. Specifically, we enforce the model to generate outputs within a structured think-answer format, compelling it to follow a "Look before Transcription" procedure. Within the <think> block, the model is required to first perform OCR to identify textual information from the slides. Subsequently, in the <answer> block, it must generate the final transcription by referring to the content in <think>, particularly using the entities identified from slides as anchors. Our training is guided by four distinct reward functions: a) Format Reward to enforce compliance with the structured output format; b) OCR Reward to promote precise extraction of textual information from slides; c) ASR Reward aimed at enhancing the overall ASR performance; and d) Visual Anchoring Reward that encourages the model to utilize the entities identified within the <think> block as references during the generation of the final transcription. Fig 1 shows an example illustrating the performance of our method. In a challenging real-world example, both a naive OLLM and a pipeline-based method fail to correctly transcribe all domain-specific entities. In contrast, our VAPO-7B model first identifies all entities in the <think> block and subsequently generates a accurate transcription, showcasing the practical advantage of our approach.
Existing SlideASR datasets, such as SlideSpeech Wang et al. (2024b) and ChineseLips Zhao et al. (2025) lack a sufficient number of specialized named entities. To address this, we constructed SlideASR-Bench, a benchmark for the entity-rich SlideASR task, which comprises two subsets: SlideASR-S and SlideASR-R. SlideASR-S is built by synthesizing slides based on entity information from the ContextASR-Bench Wang et al. (2025) dataset. To evaluate model performance in real-world scenarios, we additionally curated SlideASR-R by manually collecting 60 challenging samples from authentic professional reports across four domains: Chemistry, Medicine, Biology, and Artificial Intelligence. See Sec 5 for details.
We conduct experiments on SlideSpeech Wang et al. (2024b), ChineseLips Zhao et al. (2025) and SlideASR-Bench. Results show that our proposed approach outperforms the state-of-the-art models, e.g., Qwen3-Omni-30B-A3B Xu et al. (2025b), particularly on entity-related metrics, while maintaining accuracy on non-entity text. Ablation studies show that each reward function in VAPO is essential for achieving optimal performanc. Attention visualization shows that the VAPO model follows the "Look before Transcription" procedure. The contributions can be summarized as follows:
  • To the best of our knowledge, this work is the first to identify and analyze the limitations of OLLMs when applied to the SlideASR task.
  • We introduce VAPO, a novel post-training method designed to improve the performance of OLLMs in the SlideASR task by enforcing a structured "Look before Transcription" reasoning process.
  • We construct SlideASR-Bench, which consists of two dedicated datasets, SlideASR-S and SlideASR-R, to address the scarcity of domain-specific entities and provide robust benchmarks for the SlideASR task. The data and code will be released.
2 Related Works
Contextual ASR. The objective of contextual ASR is to incorporate contextual information, including domain labels, entity lists, and conversational history, into the speech recognition system in order to improve the recognition accuracy of named entities, and domain-specific terminology Bai et al. (2024); Xiao et al. (2025); Zhou and Li (2025). Besides textual information, researchers have focused on leveraging visual information to enhance the performance of ASR models. For example, integrating lip movement information during the recognition process Ma et al. (2023); Rouditchenko et al. (2024); Shi et al. (2022). This study focuses on the SlideASR task Zhao et al. (2025); Wang et al. (2024b, a), which involves utilizing slide content as contextual information to support the model, given that slides in presentation scenarios generally contain information closely related to the spoken content. Most existing methods for SlideASR are based on the pipeline paradigm Wang et al. (2024a); Zhao et al. (2025); Wang et al. (2024b); Yu et al. (2024); Yang et al. (2024), which results in relatively complex systems. The objective of this study is to accomplish the task using an end-to-end approach.
Omni-modal Large Language Models. Recently, OLLMs Fu et al. (2025); Yao et al. (2024); Xu et al. (2025a); Li et al. (2025a); Hu et al. (2025); Li et al. (2025b) have emerged, integrating vision, audio, and text by aligning their encoders during training for end-to-end processing. Models such as MiniCPM-o Yao et al. (2024) and Qwen2.5-Omni Xu et al. (2025a) have demonstrated strong multimodal performance. Benefiting from the unified modeling capability across visual and audio modalities, they are expected to solve the SlideASR task in an end-to-end manner. However, in practical scenarios, the models exhibit unstable behavior, for instance, they sometimes reproduce the textual content from the slides instead of generating the expected speech transcription.
Chain-of-Thought Reasoning. CoT reasoning is a breakthrough approach that enhances the complex reasoning capabilities of LLMs. Recent works Jaech et al. (2024); Guo et al. (2025) have shown that by using reinforcement learning algorithms Shao et al. (2024b); Rafailov et al. (2023); Schulman et al. (2017) to encourage models to generate intermediate reasoning steps before producing the final answer, performance on tasks involving arithmetic, commonsense, and symbolic reasoning can be significantly enhanced. This paradigm has also been extended to the multimodal domain Shao et al. (2024a); Xu et al. (2024); Ma et al. (2025); Diao et al. (2025), demonstrating the general effectiveness of making reasoning processes explicit.
3 The Failure of OLLMs in SlideASR
3.1 Problem Formulation
Figure 2: Comparison of OLLM outputs with and without slide context.
We formally define the SlideASR task as follows. Given a speech signal $A$ and a slide image $I$ synchronized with the speech, our goal is to train a model $f_{\theta}$ that takes $A$ and $I$ as joint inputs and generates the most likely corresponding text transcription $Y=\{y_{1},y_{2},…,y_{n}\}$, where $y_{n}$ is the $n$-th token in the text sequence. The objective of this task can be expressed as maximizing the conditional probability $P(Y|A,I;\theta)$, where $\theta$ represents the learnable parameters of the model.
$\theta^{*}=\arg\max_{\theta}P(Y|A,I;\theta)$ (1)
In our work, $f_{\theta}$ is an OLLM is that takes both image and audio as input and generates the final text $Y$ in an auto-regressive manner.
3.2 Observations of OLLMs on SlideASR
Although OLLMs can process both auditory and visual modalities simultaneously, we observe that their behavior on the SlideASR task is unstable. Fig 2 shows an example. We selected a sample from SlideSpeech Wang et al. (2024b) and had Qwen2.5-Omni-7B transcribe the speech under two conditions: audio-only and audio with slide image. When the model receives only audio, it produces correct ASR results. However, when both audio and the corresponding slide image are provided, the model ignores the audio signal and instead outputs the text content from the slide image, behaving like a OCR system.
To quantitatively demonstrate the instability of OLLMs’ behavior, we calculated the proportion of OCR-like behavior exhibited by four models Qwen2.5-Omni-7B Xu et al. (2025a), Qwen2.5-Omni-3B Xu et al. (2025a), MiniCPM-o-2.6 Yao et al. (2024), and Megrez-Omni Li et al. (2025a) on the SlideSpeech dataset. The judgment process involves three steps. 1) Identify Common Vocabulary: We first identify the words common to both the ground truth speech transcription and the slide text, denoted as $V_{\text{common}}$. 2) Isolate Slide-Only Vocabulary: We then create a set of slide-only words, $V_{\text{slide\_only}}$, by removing the common words from the slide’s full vocabulary. These words are only accessible to the model via OCR. 3) Detect OCR Behavior: Finally, we check for any intersection between the model’s output and $V_{\text{slide\_only}}$. If such an intersection exists, it indicates that the model has exhibited OCR behavior.
Table 1: Percentage of samples with OCR behavior on SlideSpeech dataset.
Dev set Test set
Num=1,801 Num=3,053
MiniCPM-o-2.6 57.96% 63.28%
Megrez-Omni 45.14% 44.90%
Qwen2.5-Omni-3B 15.43% 16.54%
Qwen2.5-Omni-7B 13.71% 12.87%
The results are shown in Table 1. All models exhibited a significant proportion of OCR behavior on the dataset, generating words present on the slide but absent in the audio. This fundamental and widespread failure demonstrates that simply prompting OLLMs is insufficient and that a new, more structured approach is required to control their reasoning process.
4 Visually-Anchored Policy Optimization
To address the issue of uncontrollable behavior in OLLMs for the SlideASR task, we propose Visually-Anchored Policy Optimization (VAPO). The core idea of VAPO is to transform the model from an unreliable "black-box" into a controllable one that follows a structured and effective reasoning path. This is achieved through explicit structured reasoning and multi-objective reward-based policy optimization. The process of VAPO is shown in Fig 3.
Figure 3: Overview of the Visually-Anchored Policy Optimization framework. The OLLM takes audio, slide, and instruction as input, generates a structured output, and is optimized via reward functions that guide the policy update.
4.1 Structured Reasoning Format
Inspired by CoT Reasoning Jaech et al. (2024), we design a mandatory output format to guide the model in following the reasoning principle we call "Look before Transcription". Specifically, the model is required to generate its output within a unified structure of <think><answer>.
In the <think> block, the model first processes and digests the visual information. Before generating the final transcription, it must output recognized text from the slide image $I$. This step acts as an explicit internal OCR, ensuring the model sees and understands key slide information before processing the audio. The <answer> block is where the final transcription $Y$ is generated. Instead of generating directly, the model is required to reference the visual content anchored in the <think> block. This allows specialized terms and technical jargon from the slide to serve as prior knowledge, helping the model resolve ambiguous audio and homophones, improving the transcription of key named entities. Through this format, we decompose the complex and implicit SlideASR task into an ordered, two-stage explicit reasoning process.
4.2 Reinforcement Learning Optimization
We design four reward functions to guide the model’s learning comprehensively from different dimensions. To optimize the model’s policy with these rewards, we employ the Generative Representational Policy Optimization (GRPO) Shao et al. (2024b) algorithm for fine-tuning the model.
Format Reward. This reward aims to ensure that the model’s output strictly adheres to the <think></think><answer></answer> format. A positive reward is given if the model generates the complete structure. The reward function is as follows:
$R_{\text{Format}}=\begin{cases}1,&\text{If the format is correct}\\ 0,&\text{otherwise.}\end{cases}$ (2)
OCR Reward. This reward is used to evaluate the accuracy of the content in the <think> block. We compare the text generated by the model in <think> with the actual text on the slide. We calculate the reward using the Word Error Rate (WER). If the sample is in Chinese, each character is treated as a word. To ensure the reward is non-negative, we apply a clipping mechanism that prevents it from going below zero. We denote the text in the <think> block as $T_{t}$, and the text on the slide as $T_{s}$. The reward function is as follows:
$R_{\text{OCR}}=max(1-\text{WER}(T_{t},T_{s}),0)$ (3)
ASR Reward. This reward is used to evaluate the overall quality of the final transcription text in the <answer> block. We compare the output in <answer> with the ground truth speech transcription. Similarly, we apply the same clipping mechanism in ASR reward. We denote the text in the <answer> as $T_{a}$ block, and the ground truth speech transcription as $T_{g}$. The reward function is as follows:
$R_{\text{ASR}}=max(1-\text{WER}(T_{a},T_{g}),0)$ (4)
Visual Anchoring Reward. This reward in the VAPO is responsible for establishing a connection between the <think> and <answer> blocks. This reward specifically incentivizes the model to correctly use the named entities recognized in the <think> block within the <answer> block. The calculation process is as follows. First, we extract the set of correctly identified entities $E_{\text{think}}$ from the output of the <think>. Then, we calculate the F1 score of this entity set $E_{\text{think}}$ in the output of the <answer>. The reward function is as follows:
$R_{\text{VA}}=F1_{score}(E_{think},T_{answer}).$ (5)
This reward encourages the model to reference and utilize the slide information, effectively mitigating the issue of disjoint processing between the "seeing" and "hearing".
Finally, the model’s total reward is defined as a weighted sum of the four rewards described above:
$\begin{split}R_{total}=&\lambda_{1}R_{\text{Format}}+\lambda_{2}R_{\text{OCR}}\\ &+\lambda_{3}R_{\text{ASR}}+\lambda_{4}R_{\text{VA}}\end{split}$ (6)
$\lambda$ denotes the weighting hyperparameters for each individual reward.
5 SlideASR-Bench: A Benchmark for entity-rich SlideASR Task
Our goal is to improve the accuracy of model transcription for domain-specific entities in the context of slides. Existing public SlideASR datasets, such as SlideSpeech Wang et al. (2024b) and ChineseLips Zhao et al. (2025), provide a valuable benchmark for real-world general-domain scenarios. However, we observe that these datasets often lack a sufficient density of domain-specific named entities in both speech and slides. This creates a significant challenge in training a model capable of handling entity-rich slides and speech, as well as objectively evaluating its performance in entity-rich SlideASR scenarios. To address this critical resource bottleneck, we constructed SlideASR-Bench, which includes two datasets: a synthetic dataset, SlideASR-S, for training and evaluation, and a challenging evaluation set, SlideASR-R, for assessing real-world performance. Table 2 presents the detailed information.
SlideASR-S. To train and evaluate models for entity-rich scenarios, we constructed SlideASR-S based on the ContextASR-Bench Wang et al. (2025) dataset. ContextASR-Bench leverages LLMs, such as DeepSeek-R1 Guo et al. (2025), to generate colloquial text rich in named entities based on seed text. The seed text is sourced from Named Entity Recognition (NER) datasets across multiple domains, such as medicine, culture, ecology. Then, models such as CosyVoice2 Du et al. (2024) and XTTS-v2 Casanova et al. (2024), which are text-to-speech models, are used to convert the generated text into natural and fluent speech.
We extract metadata for each sample from ContextASR-Bench, including the domain label $L_{\text{domain}}$ and the list of domain-specific entities $E=\{e_{1},e_{2},\dots,e_{k}\}$. Using the $L_{\text{domain}}$ and $E$ as input, we employ a LLM (e.g., Qwen2.5-14B-Instruct  [^1] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct) to generate a short descriptive text in slide style. The prompt guides the LLM to include a title and key points, ensuring that all entities from the original audio are naturally embedded in the generated text. The prompt can be found in Appendix  A. Finally, we use Python’s Matplotlib library to render the text into an image, simulating a presentation slide. Each generated image serves as synthetic visual context.
SlideASR-R. Although synthetic data is useful for model training and evaluation, there remains a certain domain gap between synthetic and real-world scenarios. To assess the model’s generalization ability in real, complex environments, we manually constructed a small-scale, high-quality, and challenging test set.
We collected 60 real presentation audio clips and corresponding slide images from publicly available academic report videos, covering four specialized domains: chemistry, medicine, biology, and artificial intelligence. For each sample, we manually annotated the data by carefully comparing the speech and slide image, identifying the domain-specific entities that appear in both. We named this dataset SlideASR-R (R for Real), which contains 200 domain-specific entities from real-world scenarios. Despite its relatively small size, it provides a highly challenging benchmark for assessing model performance in practical applications.
Table 2: Details of our proposed SlideASR-Bench.
#Sample #Entity #Hour
SlideASR-S (Train set) 6,413 44,240 67.3
SlideASR-S (Test set) 2,054 13,895 18.5
SlideASR-R 60 200 0.35
6 Experiment
6.1 Setup
Table 3: Results on the SlideSpeech, a real-world English SlideASR dataset. $\dagger$ represents results from the original paper. The best and second-best results are in bold and underlined, respectively.
Model Dev Test
[…]
Contextless
Qwen2-Audio 12.56 12.85 8.72 91.43 13.19 13.59 7.53 92.91
Mi-Dasheng 13.63 13.97 9.06 90.97 14.61 15.04 8.52 91.59
MiniCPM-o-2.6 16.09 16.68 8.14 91.98 18.71 19.41 8.90 91.50
Qwen2.5-Omni-3B 15.53 16.22 6.30 93.76 12.00 12.45 5.72 94.41
Qwen2.5-Omni-7B 11.75 12.20 5.39 94.78 11.75 12.20 5.39 94.78
Qwen3-Omni-30B-A3B 10.87 11.31 5.02 95.04 11.71 12.21 4.64 95.50
Slide text as context (Pipeline)
Qwen2-Audio 139.81 145.05 69.94 85.40 146.08 152.41 56.99 88.98
Mi-Dasheng 33.67 35.18 13.56 93.02 47.21 49.34 17.21 91.00
Qwen3-Omni-30B-A3B 50.43 52.85 18.05 96.45 57.12 59.27 26.75 96.34
[…]
Slide image as context (End-to-End)
MiniCPM-o-2.6 182.96 192.83 51.07 86.26 210.37 220.96 60.92 83.22
Qwen2.5-Omni-3B 12.22 12.74 5.26 95.17 19.99 20.71 9.80 94.44
Qwen2.5-Omni-7B 13.65 14.13 7.19 92.84 14.97 15.58 6.33 93.99
Qwen3-Omni-30B-A3B 19.85 20.64 9.30 95.59 24.13 24.88 13.44 94.74
VAPO-3B (Ours) 9.84 10.31 3.61 96.54 10.73 11.24 3.55 96.57
VAPO-7B (Ours) 8.62 9.08 2.48 97.61 10.31 10.84 2.87 97.32
Table 4: Results on the SlideASR-Bench. The best and second-best results are in bold and underlined, respectively.
Model SlideASR-S (en) SlideASR-S (zh) SlideASR-R
WER NE-WER NE-FNR WER NE-WER NE-FNR NE-WER NE-FNR
Contextless
Qwen2-Audio 11.90 36.29 47.84 6.02 22.83 40.36 74.56 76.73
Mi-Dasheng 12.16 30.32 35.40 4.67 19.33 35.81 54.08 61.39
MiniCPM-o-2.6 11.19 27.51 30.93 10.35 25.00 41.62 55.85 65.37
Qwen2.5-Omni-3B 8.37 24.15 31.04 4.47 19.89 38.08 61.31 66.83
Qwen2.5-Omni-7B 8.15 23.44 27.77 4.34 17.54 32.80 53.68 63.37
Qwen3-Omni-30B-A3B 9.06 14.61 15.53 20.77 23.31 22.49 40.43 41.09
Slide text as context (Pipeline)
Qwen2-Audio 92.16 66.38 24.82 39.09 50.58 31.52 59.04 21.29
Mi-Dasheng 78.98 49.85 30.58 66.88 56.30 32.30 47.52 26.73
Qwen3-Omni-30B-A3B 34.65 32.35 8.56 9.76 15.85 13.54 34.01 28.22
Slide image as context (End-to-End)
MiniCPM-o-2.6 112.90 49.65 15.01 89.53 61.25 45.67 63.73 66.83
Qwen2.5-Omni-3B 100.08 53.19 18.72 86.86 65.62 9.62 49.00 53.47
Qwen2.5-Omni-7B 57.21 35.76 15.04 91.83 54.04 3.36 41.77 35.15
Qwen3-Omni-30B-A3B 101.45 59.64 12.08 79.21 46.45 5.54 32.26 24.75
VAPO-3B (Ours) 4.90 3.19 3.73 2.47 4.21 2.22 27.28 19.31
VAPO-7B (Ours) 4.60 2.83 2.97 2.13 3.78 1.36 26.48 15.35
Figure 4: Attention visualization. Left: input image and transcribed audio text. Right: attention flows.
We fine-tune the Qwen2.5-Omni-3B and Qwen2.5-Omni-7B models on the SlideASR-S training set using the proposed VAPO algorithm. All experiments were run on 4 NVIDIA A100 GPUs with 80 GB of memory each. The weights of the reward functions, $\lambda_{1}$ to $\lambda_{4}$, are all set to 1.
We evaluate models on SlideSpeech Wang et al. (2024b) and SlideASR-Bench, using three settings: Contextless, where only audio is used; Slide text as context, a pipeline setting where OCR extracts text from slides and combines it with audio; and Slide image as context, an end-to-end setting that directly inputs both slide images and audio. We select mainstream LALMs Chu et al. (2024); Dinkel et al. (2025) and OLLMs Yao et al. (2024); Xu et al. (2025a, b) as baselines. Note that the same model may support multiple settings. For example, Qwen3-Omni-30B-A3B Xu et al. (2025b) can take audio alone as input or accept text or image as context. Following Zhao et al. (2025), we use PaddleOCR  [^2] https://github.com/PaddlePaddle/PaddleOCR. for OCR. See Appendix B for details of evaluation prompts, metrics and baseline models. Additionally, we present the results on ChineseLips Zhao et al. (2025), a real-world Chinese dataset, in Appendix C (Table 6) to demonstrate VAPO’s generalization in real-world scenarios.
6.2 Main Results
Results on SlideSpeech. Table 3 shows the results on the real-world SlideSpeech dataset. It can be observed that baseline models generally perform poorly when incorporating slide information, with performance even degrading compared to the contextless setting. For instance, the Qwen3-Omni-30B-A3B Xu et al. (2025b), using either slide text or the slide image as context exhibits a higher WER compared to using audio alone. In contrast, our proposed VAPO method achieves the best results on the SlideSpeech dataset. The VAPO-7B model reaches a WER of 10.31 and a recall of 97.32 on the test set, outperforming baselines such as Qwen3-Omni-30B-A3B (Recall=95.5) and MaLa-ASR Yang et al. (2024) (WER=11.26).
Results on SlideASR-Bench. Table 4 presents the results on SlideASR-Bench. Unlike SlideSpeech, on SlideASR-Bench, all audio-only models struggle to recognize specialized named entities. For example, on the SlideASR-R dataset, even the strongest audio-only model, Qwen3-Omni-30B-A3B, reaches a NE-FNR of 41.09. This underscores the essential role of incorporating visual information in entity-rich ASR tasks. Additionally, baseline models fail to effectively utilize contextual information, whether in text or image form. For example, on SlideASR-S, when Qwen3-Omni-30B-A3B uses OCR text or slide images as context, although NE-WER decreases compared to audio-only input, the overall WER increases significantly. This aligns with our observations in Section 3, the model is not transcribing the speech properly but rather outputting the context content instead.
The VAPO method achieves significant improvements in recognizing key entities. For example, on the challenging real-world evaluation set SlideASR-R, VAPO-7B reduces the NE-FNR from the baseline best of 28.22 to 15.35. These results indicate that VAPO can accurately extract and anchor entity text from slides, thereby improving transcription accuracy. See Appendix D, for the successful cases and analysis of a failure case.
6.3 Ablation Results
Table 5: Ablation results on the SlideASR-R dataset.
ASR
Reward
OCR
Reward
VA
Reward
[…]
Qwen2.5-Omni-3B
49.00 53.47
37.23 31.19
29.97 22.28
27.28 19.31
Qwen2.5-Omni-7B
41.77 35.15
28.63 20.30
26.75 18.32
26.48 15.35
Table 5 presents the ablation results for different reward functions. The results show that, the ASR reward enables the models to initially reference slide information. The OCR reward further improves the accuracy of slide text extraction, enhancing overall performance. Finally, adding the Visual Anchoring (VA) reward strengthens the models’ focus on key entities in <think>, achieving the best results. Additional ablation results for SlideSpeech (Table 7) and SlideASR-S (Table 8) are in Appendix E.1. Table 9 in Appendix E.2 presents the ablation results for the reward weights ($\lambda_{1}$$\lambda_{2}$$\lambda_{3}$$\lambda_{4}$), demonstrating the trade-off between overall transcription accuracy and entity recognition performance, and how different weight configurations impact this balance. The 1:1:1:1 configuration optimally balances both.
6.4 Attention Visualization
To examine the model’s behavior during final transcription, we visualize the attention weights of VAPO-7B on a SlideSpeech case, as shown in Fig 4. We observed that when the model generates speech transcription in <answer>, it refers to the entity information in <think>. For instance, for the key entity "Concentriq", after generating the token "Concent" <answer>, the model pays significant attention to the "ri" token in <think> and subsequently generates it. It then refers to the "q" token in the <think> block, enabling accurate and complete transcription of "Concentriq". A similar process occurs for the entity "proscia". This demonstrates that the model indeed references the slide information during transcription. Further cases can be found in Appendix F.
7 Conclusion
This paper identifies key failures in OLLMs when applied to the SlideASR task, where they often disregard audio input and function merely as OCR systems. To address this, we introduce Visually-Anchored Policy Optimization (VAPO), a novel training method that enforces a structured "Look before Transcription" reasoning process. By leveraging a <think><answer> format and multi-faceted reward functions, VAPO significantly enhances the model’s ability to integrate visual and auditory information. Furthermore, we developed SlideASR-Bench, an entity-rich benchmark, to facilitate more robust training and evaluation. Extensive experiments demonstrate that VAPO substantially improves the transcription accuracy, particularly for specialized terms, establishing a more effective end-to-end paradigm for SlideASR.
8 Limitations
While VAPO demonstrates significant improvements, this work has several limitations.
Task Generalization. Our current approach is highly specialized for leveraging textual information from presentation slides and does not incorporate other visual cues, such as images of entities (e.g., pictures of specific drugs). In the future, we will adapt our "Look before Transcription" paradigm to handle more diverse multimodal environments, where various visual elements play a crucial role.
Real-World Robustness. While our training relies on the synthetic SlideASR-S dataset, we have validated its effectiveness on three real-world datasets, SlideSpeech, ChineseLips and SlideASR-R. Nonetheless, a subtle domain gap may still exist, as synthetic slides may not fully capture the stylistic diversity and visual noise (e.g., complex diagrams, low-quality images) of all real-world presentations.
Inference Efficiency: The structured reasoning process of VAPO introduces a computational overhead, resulting in higher inference latency compared to models of the same size (detailed in Appendix G). This makes VAPO most suitable for offline applications where accuracy is critical. However, this trade-off between latency and accuracy is often acceptable for offline transcription tasks where precision is paramount. We will explore strategies such as model distillation in our future research to improve efficiency, enabling its use in real-time applications.
9 Ethical Considerations
The primary societal benefit of our work is enhancing accessibility by improving the transcription accuracy of specialized terms, which can significantly aid individuals who are deaf or hard of hearing. However, this technology must be deployed with caution in high-stakes settings, such as medical transcription, where errors could lead to serious consequences. We advocate for responsible development and believe that human oversight is essential for any critical applications.
Appendix
Appendix A Prompts for Generate SlideASR-S
The prompt for the LLM to generate text paragraphs based on a domain label and an entity list is as follows.
Prompt for Qwen2-14B-Instruct to generate slide text Given a domain label and a list of entities, generate a title and a paragraph for use in a PPT report, with the requirement that the paragraph includes these entities, Keep paragraphs within 150 words.
Domain label:
List of entities:
Output format:
Title
Paragraph
Appendix B Evaluation Details
B.1 Prompts
The prompts for baseline models and our VAPO models are as follows.
Prompt for baseline models ### Contextless
Convert the audio to text.
### Slide text as context
The speech is the speaker’s talk accompanied by a slide, with the text of the slide being: {}
Transcribe the speech into text by integrating the speech with the slide content.
### Slide image as context
Taking the image content into account, convert the audio to text.
Prompt for VAPO model ### Slide image as context
Role:System
Your task is to convert the speech into text, and the image serves as the reference content related to the speech.
Role:User
First, recognize the text in the image and output it within <think> </think>. Then, referring to the thinking content, output the speech recognition result within <answer> </answer>
B.2 Metrics
For SlideSpeech, as in the original work Wang et al. (2024b), we use four metrics
  • WER: word error rate.
  • U-WER: unbiased word error rate, computed on non-keyword segments, to evaluate model impact on general transcription.
  • B-WER: unbiased word error rate, which measures errors on keyword spans.
  • Recall: keyword recall, the percentage of keywords fully and correctly recognized.
For SlideASR-Bench, we maintain consistency with ContextASR-Bench Wang et al. (2025) and use the following three evaluation metrics:
  • WER: word error rate. For Chinese samples, we treat each character as a word.
  • NE-WER: WER of named entity portion, we first perform a fuzzy match to identify key entities (with an edit distance tolerance of $\frac{2}{WordCountOfEntity}-1$) in the model’s output, and then calculate the WER based on the fuzzy-matched entities.
  • NE-FNR: The false negative ratio of named entities, calculated as $1-\frac{N}{H}$, where $H$ and $N$ denote the recognized and ground-truth entity counts.
Table 6: Results on ChineseLips, a real-world Chinese SlideASR dataset. The best and second-best results are in bold and underlined, respectively.
[…]
Contextless
Qwen2-Audio 12.536
Mi-Dasheng 3.311
MiniCPM-o-2.6 2.252
Qwen2.5-Omni-3B 1.937
Qwen2.5-Omni-7B 2.243
Qwen3-Omni-30B-A3B 2.202
Slide text as context (Pipeline)
Qwen2-Audio 84.291
Mi-Dasheng 65.505
Qwen3-Omni-30B-A3B 69.172
Slide image as context (End-to-End)
MiniCPM-o-2.6 76.203
Qwen2.5-Omni-3B 24.847
Qwen2.5-Omni-7B 14.340
Qwen3-Omni-30B-A3B 41.930
VAPO-3B (Ours) 1.548
VAPO-7B (Ours) 1.298
B.3 Baselines
LALMs. For LALMs, we selected Qwen2-Audio Chu et al. (2024) and Mi-Dasheng Dinkel et al. (2025) as baselines, both with 7B parameters. ASR is a core capability of these models. Additionally, they have instruction-following abilities, making them suitable for the context-enhanced ASR task, i.e., Slide text as context setting. For SlideSpeech, we additionally selected LCB-net Yu et al. (2024) and MaLa-ASR Yang et al. (2024) as baselines. These models were trained on the SlideSpeech training set, and the results for Dev and Test sets are provided Yang et al. (2024).
OLLMs. For OLLMs, we selected MiniCPM-o-2.6 Yao et al. (2024), Qwen2.5-Omni-3B Xu et al. (2025a), Qwen2.5-Omni-7B Xu et al. (2025a) and Qwen3-Omni-30B-3B Xu et al. (2025b) as baselines. Similarly, these models not only have ASR capabilities and instruction-following abilities, but they can also directly accept both image and audio as inputs.
Appendix C Results on ChineseLips
Table 6 presents the results on ChineseLips Zhao et al. (2025). Similar to SlideSpeech Wang et al. (2024b), ChineseLips is a real-world general-domain SlideASR dataset with low entity density both in the speech and the slides. Since ChineseLips does not provide text information for the slides, we report the CER metric on the transcribed text.
The results show that our method achieves the lowest CER, demonstrating its effectiveness in real-world general-domain scenarios.
Figure 5: Qualitative examples from SlideASR-R.
Appendix D Case Study
Figure 6: A failure case from SlideASR-R. OCR errors occurred due to low image resolution and small entity font size, leading to the loss of correct visual anchor points.
D.1 Successful Case
Table 7: Ablation results on the SlideSpeech dataset.
ASR OCR VA Dev Test
[…]
12.22 12.74 5.26 95.17 19.99 20.71 9.80 94.44
10.22 10.68 4.01 96.08 11.02 11.52 3.92 96.17
9.97 10.49 3.83 96.30 11.74 12.25 4.47 95.63
9.84 10.31 3.61 96.54 10.73 11.24 3.55 96.57
Table 8: Ablation results on the SlideASR-S dataset.
ASR OCR VA En Zh
[…]
100.08 53.19 18.72 86.86 65.62 9.62
5.40 3.78 4.47 2.49 4.33 2.49
4.98 3.71 4.23 2.58 4.54 2.82
4.90 3.19 3.73 2.47 4.21 2.22
Fig 5 shows a comparison of outputs from Qwen2.5-Omni-7B, and Qwen3-Omni-30B-A3B and our proposed VAPO-7B models on samples from the SlideASR-R dataset. Among them, Qwen2.5-Omni-7B uses audio-only input, Qwen3-Omni-30B-A3B uses OCR text extracted from the slide image as context, and VAPO-7B uses the slide image as context input. For Qwen2.5-Omni-7B, due to the lack of auxiliary information, the entity error rate is relatively high. For example, it misrecognized "ConVIRT" as "convert". For Qwen3-Omni-30B-A3B, although slide text is used as context, it fails to utilize it effectively. For example, it also misrecognized "ConVIRT" as "Convert". The VAPO-7B model achieves higher entity recognition accuracy thanks to its "Look before Transcription" reasoning structure.
D.2 Failure Case
Fig 6 reveals a failure mode of our VAPO-7B model, originating from a visual perception error. The issue begins in the <think> block, where the internal OCR component misidentifies a key entity. Specifically, the correct character (pronounced ke) is misrecognized as a different character (pronounced pin). This OCR error is highly plausible because the two characters are visually confusable due to their structural similarity. This type of resemblance is a known challenge for OCR systems, especially with low-resolution text, and it results in the absence of a correct visual anchor.
Due to the lack of a correct visual anchor in the <think> block, the entity transcription in the <answer> block ends up being incorrect. This case highlights that VAPO’s performance is sensitive to low resolution and font size, particularly with visually similar characters. However, this doesn’t diminish its overall advantage, as a contextless audio-only model would be equally, if not more, prone to failure when confronted with such inherent ambiguities in the source modalities.
Table 9: Ablation results of reward function weights on SlideASR-Bench.
Hyperparameter SlideASR-S (en) SlideASR-S (zh) SlideASR-R
$\lambda_{1}:\lambda_{2}:\lambda_{3}:\lambda_{4}$ WER NE-WER NE-FNR WER NE-WER NE-FNR WER NE-FNR
$1:1:1:1$ 4.90 3.19 3.73 2.47 4.21 2.22 27.28 19.31
$1:1:1:2$ 5.27 3.34 3.78 2.50 4.30 2.09 27.67 17.73
$1:1:2:1$ 5.32 4.12 3.91 2.48 4.38 2.09 30.35 22.17
$1:2:1:1$ 5.17 3.45 3.80 2.51 4.23 1.99 27.54 21.18
Table 10: Comparison of inference time and NE-FNR on the SlideASR-R dataset.
Model Setting
Inference time per sample (s)
NE-FNR
Qwen3-Omni-30B-A3B Slide text as context 105.98 28.22
Qwen3-Omni-30B-A3B Slide image as context 172.95 24.75
Qwen2.5-Omni-7B Slide image as context 2.51 35.15
VAPO-7B Slide image as context 7.27 15.35
Appendix E More Ablation Results
E.1 Ablation Results of Reward Components
Table 7 and Table 8 respectively present the ablation results of VAPO-3B (based on Qwen2.5-Omni-3B) on SlideSpeech Wang et al. (2024b) and SlideASR-S. The results indicate that different reward functions have a positive impact on the final performance.
Figure 7: More cases of attention visualization.
E.2 Sensitivity Analysis of Reward Weights
To investigate the sensitivity of our VAPO method to the weights ($\lambda_{1}:\lambda_{2}:\lambda_{3}:\lambda_{4}$) of the four reward functions, we conducted an ablation study on the SlideASR-Bench. As shown in Table 9, we compared our default equal weighting scheme (1:1:1:1) against configurations where the weight of the OCR reward ($\lambda_{2}$), ASR reward ($\lambda_{3}$), or the Visual Anchoring reward ($\lambda_{4}$) was doubled.
The results confirm that the 1:1:1:1 configuration provides the best overall performance, achieving the lowest Word Error Rate (WER) across all three subsets. This demonstrates the importance of a balanced approach to synergistic learning.
The analysis also reveals a key trade-off. While doubling the Visual Anchoring reward (1:1:1:2) achieves the best NE-FNR (17.73) on the challenging real-world SlideASR-R set, it does so at the cost of higher overall WER. Conversely, overweighting the ASR reward (1:1:2:1) proved counterproductive, as it discourages the model from leveraging crucial visual context and significantly degrades performance.
Therefore, we conclude that an equal weighting of the reward functions is a simple and effective choice for the VAPO method. This configuration was used for all other experiments reported in this paper.
Appendix F More Cases of Attention Visualization
Figure 6 further presents two cases of attention visualization. It can be seen that when transcribing key entities, the model is able to focus its attention on the same entities in the <think> block. This desirable property enables the model to accurately transcribe key entities in the speech.
Appendix G Inference Latency Analysis
To evaluate the practical inference efficiency of our proposed VAPO framework, we measured the average inference time per sample on SlideASR-R and compared it against several key baseline models. The results are presented in Table 10.
As shown, our VAPO-7B model has an inference time of 7.27 seconds per sample. This is slower than Qwen2.5-Omni-7B (2.51s), which is expected, as the structured <think><answer> generation process introduces a computational overhead. However, this moderate increase in latency is accompanied by a dramatic improvement in accuracy, with the NE-FNR dropping from 35.15 to 15.35.
More importantly, when compared to the best baseline model Qwen3-Omni-30B-A3B, our VAPO-7B is significantly more efficient and accurate. While not yet suitable for real-time applications, the latency of VAPO is a reasonable trade-off for its state-of-the-art performance, particularly for offline transcription tasks where accuracy is paramount.

### v2

@@
-Look before Transcription: End-to-End SlideASR with Visually-Anchored Policy Optimization
+VAPO: End-to-end Slide-Enhanced Speech Recognition with Omni-modal Large Language Models
 Abstract
-Automatic speech recognition (ASR) systems often struggle with domain-specific terminology, especially in specialized settings such as academic lectures.
-To address this, we define the SlideASR task, which leverages the rich visual information from presentation slides to improve transcription accuracy.
-Existing pipeline methods for this task tend to be complex and underperform.
-Although omni-modal large language models (OLLMs) provide a promising end-to-end framework, they frequently fail in practice by degenerating into simple optical character recognition (OCR) systems.
-To overcome this, we propose Visually-Anchored Policy Optimization (VAPO), a novel post-training method designed to control the model’s reasoning process.
-Drawing on the Chain-of-Thought reasoning paradigm, VAPO enforces a structured "Look before Transcription" procedure using a <think><answer> format.
-Specifically, the model first performs OCR on the slide content within the think step, then generates the transcription by referencing this recognized visual information in the answer step.
-This reasoning process is optimized via reinforcement learning with four distinct rewards targeting format compliance, OCR accuracy, ASR quality, and visual anchoring consistency.
-To support further research, we construct SlideASR-Bench, a new entity-rich benchmark consisting of a synthetic dataset for training and testing, and a challenging real-world set for evaluation.
-Extensive experiments demonstrate that VAPO significantly improves recognition of domain-specific terms, establishing an effective end-to-end paradigm for SlideASR.
-Look before Transcription: End-to-End SlideASR with
-Visually-Anchored Policy Optimization
-[…]
+Omni-modal large language models (OLLMs) offer a promising end-to-end solution for slide-enhanced speech recognition due to their inherent multimodal capabilities.
+However, we found a fundamental issue faced by OLLMs: Visual Interference, where models show a bias towards visible text over auditory signals, causing them to hallucinate slide content that was never spoken.
+To address this, we propose Visually-Anchored Policy Optimization (VAPO), which aims to reshape models’ inference process to follow the human-like “Look-then-Listen” inference chain.
+Specifically, we design a temporally decoupled policy: the model first extracts visual priors in a <think> block to serve as semantic anchors, then generates the transcription in an <answer> block.
+The policy is optimized via multi-objective reinforcement learning.
+Furthermore, we introduce SlideASR-Bench, a comprehensive benchmark designed to address the scarcity of entity-rich data, comprising a large-scale synthetic corpus for training and a challenging real-world test set for evaluation.
+We conduct extensive evaluations demonstrating that VAPO effectively eliminates visual interference and achieves state-of-the-art performance on SlideASR-Bench and public datasets, significantly reducing entity recognition errors in specialized domains.
+VAPO: End-to-end Slide-Enhanced Speech Recognition with Omni-modal Large Language Models
+Rui Hu1,2, Delai Qiu2, Yining Wang2, Shengping Liu2, Jitao Sang1,3thanks: Corresponding author.
+1Beijing Key Laboratory of Traffic Data Mining and Embodied Intelligence, Beijing Jiaotong University 2Unisound AI Technology Co., Ltd.
+3State Key Laboratory of AI Safety, Beijing {rui.hu, jtsang}@bjtu.edu.cn https://github.com/isruihu/SlideASR-Bench
 1 Introduction
-Current end-to-end automatic speech recognition (ASR) models, such as Whisper Radford et al.
-(2023), have demonstrated impressive performance in transcribing common words.
-However, recognition performance often deteriorates significantly in specialized domains, such as academic lectures, technical presentations, or medical seminars.
-Previous works have improved ASR accuracy by incorporating lip movement information from the speaker Afouras et al.
+Figure 1: Comparison of paradigms for the SlideASR task.
+(a) Pipeline paradigm: Cascades independent modules which is complex.
+(b) End-to-end paradigm: Illustrates two approaches using an OLLM.
+The Vanilla path directly generates the ASR result, often leading to visual interference.
+In contrast, our Look-then-Listen path enforces a structured inference chain, explicitly decoupling visual perception from auditory processing.
+Current Automatic Speech Recognition (ASR) models, such as Whisper Radford et al.
+(2023), have demonstrated impressive performance in general domains.
+However, recognition accuracy often deteriorates significantly in specialized scenarios, such as academic lectures or technical presentations where domain-specific terminologies and rare entities are prevalent Sinhamahapatra and Niehues (2025).
+While previous works have improved ASR accuracy by incorporating lip movement information Afouras et al.
 (2022); Ma et al.
 (2021, 2023); Shi et al.
-(2022).
-However, in addition to facial information, there is multimodal textual information on the slide that is closely related to the current speech of the speaker Wang et al.
-(2024b); Zhao et al.
-(2025), which lipreading-based approaches cannot utilize.
-For clarity, we refer to the task of improving ASR accuracy by incorporating visual cues from presentation slides as SlideASR.
-Figure 1: We compare the outputs of Qwen2.5-Omni-7B, Qwen3-Omni-30B-A3B (with OCR text as context), and our VAPO-7B on a real Chinese medical report sample.
-Red text indicates incorrectly transcribed named entities.
-Currently, the dominant strategy for SlideASR task is the pipeline paradigm.
-Specifically, these methods can be divided into two categories.
-The first is post-processing correction Trinh et al.
-(2025), which utilizes large language models (LLMs) for text-level refinement after independent ASR and optical character recognition (OCR).
-The second is contextual enhancement Wang et al.
-(2024b, a); Yu et al.
-(2024); Yang et al.
-(2024), where OCR-extracted text is injected as context into the ASR model.
-The post-processing correction methods involve multiple modules, resulting in a complex workflow and relatively high latency.
-In contrast, for contextual enhancement methods, we observed that when OCR text is provided to large audio language models (LALMs) Chu et al.
+(2022), they ignore the rich semantic context present in presentation slides.
+In these scenarios, critical keywords and proper nouns are often explicitly displayed on the slides, serving as strong visual cues for the spoken content Wang et al.
+(2024a, b).
+For clarity, we refer to the task of improving ASR accuracy by incorporating visual context from presentation slides as SlideASR.
+Currently, the dominant strategy for the SlideASR task is the pipeline paradigm Wang et al.
+(2024b).
+As shown in Fig.
+1(a), this paradigm extracts text from slide images and selects domain-specific words, which are then fed as textual context into Large Audio Language Models (LALMs) Chu et al.
+(2024); Bai et al.
 (2024); Dinkel et al.
-(2025), the models often tend to repeat the context rather than effectively assisting the recognition process.
-Beyond the pipeline paradigm, we argue that the recently emerging omni-modal large language models (OLLMs) Xu et al.
+(2025) for ASR.
+This cascading process involves multiple modules, suffering from complexity and error accumulation.
+Given these limitations, a natural question arises: Can we establish an end-to-end paradigm that directly utilizes the slide image as visual context, thus bypassing the complex pipeline?
+To answer this question, we argue that the recently emerging Omni-modal Large Language Models (OLLMs) Xu et al.
 (2025a, b); Yao et al.
 (2024); Li et al.
-(2025a) are inherently well-suited for the SlideASR task.
-OLLMs are capable of simultaneously processing textual, visual, and auditory modalities, and performing tasks based on human instructions.
-Therefore, they are theoretically able to accomplish the SlideASR task in an end-to-end manner.
-However, we identify two key issues with existing OLLMs when applied to the SlideASR task.
-First, their behavior is often uncontrollable.
-When given an audio clip and the corresponding slide image, the models do not consistently perform accurate transcription.
-In some cases, they simply copy text from the slide, effectively acting as OCR systems instead of transcribing speech.
-Second, even when ASR is successful, the model often fails to leverage entity information from the slide.
-Entities that are clearly shown in the image are sometimes transcribed incorrectly, suggesting weak integration of visual context.
-These issues highlight that the implicit and opaque reasoning processes of OLLMs are unreliable for complex multimodal tasks such as SlideASR.
-Inspired by recent advances in LLMs, where explicit Chain-of-Thought (CoT) reasoning Guo et al.
-(2025); Jaech et al.
-(2024); Shao et al.
-(2024b) has been shown to improve performance on complex problems, we argue that incorporating structured reasoning into the SlideASR task can help constrain model behavior and address the challenges of multimodal fusion.
-To this end, we propose Visually-Anchored Policy Optimization (VAPO), a post-training method tailored to enhance the performance of OLLMs on the SlideASR task.
-The core idea of VAPO is to replace the model’s implicit and uncontrolled reasoning with an explicit, structured process.
-Specifically, we enforce the model to generate outputs within a structured think-answer format, compelling it to follow a "Look before Transcription" procedure.
-Within the <think> block, the model is required to first perform OCR to identify textual information from the slides.
-Subsequently, in the <answer> block, it must generate the final transcription by referring to the content in <think>, particularly using the entities identified from slides as anchors.
-Our training is guided by four distinct reward functions: a) Format Reward to enforce compliance with the structured output format; b) OCR Reward to promote precise extraction of textual information from slides; c) ASR Reward aimed at enhancing the overall ASR performance; and d) Visual Anchoring Reward that encourages the model to utilize the entities identified within the <think> block as references during the generation of the final transcription.
-Fig 1 shows an example illustrating the performance of our method.
-In a challenging real-world example, both a naive OLLM and a pipeline-based method fail to correctly transcribe all domain-specific entities.
-In contrast, our VAPO-7B model first identifies all entities in the <think> block and subsequently generates a accurate transcription, showcasing the practical advantage of our approach.
-Existing SlideASR datasets, such as SlideSpeech Wang et al.
+(2025a) offer a promising solution.
+OLLMs are capable of simultaneously processing textual, visual, and auditory modalities.
+Thus, they are inherently well-suited to accomplish the SlideASR task in an end-to-end manner.
+However, our preliminary investigation reveals a critical gap between this theoretical potential and practical performance (Sec.
+2).
+Specifically, we observe that current OLLMs suffer from severe visual interference.
+Instead of using the slide image as auxiliary information, the models exhibit a strong tendency to hallucinate, i.e., incorrectly transcribe visible slide text while this text is absent from the spoken audio.
+We attribute this failure to the lack of an intermediate reasoning phase.
+Considering human behavior, when listening to professional presentations, humans typically adopt a “Look-then-Listen” strategy, which temporally decouples the processing of the two modalities.
+We first scan the slide to establish a contextual prior of the topic and then anchor the auditory input within this context.
+In contrast, vanilla OLLMs process visual and auditory signals simultaneously (left path in Fig.
+1(b)).
+Lacking this sequential guidance, strong visual cues tend to suppress auditory inputs, resulting in the observed interference.
+Inspired by this, we propose Visually-Anchored Policy Optimization (VAPO) which centers on reshaping the model’s inference process to align with the human-like “Look-then-Listen” workflow (right path in Fig.
+1(b)).
+Specifically, we design a <think><answer> format to explicitly decompose the multimodal task.
+In the <think> phase, the model is required to first perform Optical Character Recognition (OCR) to establish a visual context prior.
+Subsequently, in the <answer> phase, it generates the transcription by attending to the audio while referencing the extracted content as a reliable anchor.
+To instantiate this policy, we employ four complementary rewards: a) Format Reward to ensure structural compliance; b) OCR Reward to promote precise visual perception; c) ASR Reward to maintain the overall transcription accuracy; and d) Visual Anchoring Reward to encourage the model to effectively leverage key entities identified within the <think> phase.
+Furthermore, to address data bottlenecks, we construct SlideASR-Bench.
+Existing datasets, e.g., SlideSpeech Wang et al.
+(2024b), primarily focus on general scenarios and lack sufficient entity density.
+SlideASR-Bench is specifically tailored to promote the utilization of visual cues for specialized terms and comprises two subsets: 1) SlideASR-S, a large-scale synthetic corpus derived from ContextASR-Bench Wang et al.
+(2025) providing both training and test sets; and 2) SlideASR-R, a small real-world test set for evaluation in complex presentation environments.
+We conduct extensive experiments on SlideSpeech Wang et al.
+(2024b) and our proposed SlideASR-Bench.
+Empirical results demonstrate that our approach significantly outperforms state-of-the-art models, e.g., Qwen3-Omni Xu et al.
+(2025b), particularly on entity-related metrics.
+The main contributions are summarized as follows:
+  • Analysis: We identify visual interference as the primary bottleneck in current OLLMs for SlideASR, revealing why naive end-to-end approaches fail.
+  • Method: We propose VAPO, which reshapes the inference process into a structured, human-like “Look-then-Listen” workflow, explicitly temporally decoupling visual perception from auditory transcription to eliminate visual interference.
+  • Data: We construct SlideASR-Bench, comprising synthetic and real-world subsets, to train and evaluate SlideASR task in entity-rich scenarios.
+2 Visual Interference of OLLMs in End-to-end SlideASR
+Although OLLMs theoretically have the capability to process auditory and visual modalities simultaneously, our preliminary investigation reveals a critical failure mode in their practical application.
+We present a motivating example in Fig.
+2 using Qwen2.5-Omni-7B Xu et al.
+(2025a).
+In the audio-only setting, the model accurately transcribes the speech.
+However, when the corresponding slide image is introduced, the model’s behavior shifts drastically: instead of utilizing the visual information as auxiliary context, it ignores the auditory signal and directly outputs the text visible on the slide.
+We term this specific failure mode as Visual Interference.
+Formally, it is defined as a phenomenon where the model generates text present in the visual context while it is absent from the speech.
+Figure 2: An illustrative example of Visual Interference in the end-to-end SlideASR task.
+Left: the model correctly transcribes the speech in audio-only mode.
+Right: the model erroneously outputs the text on the slide when the image is introduced as visual context.
+To systematically quantify this phenomenon, we introduce a metric termed the Visual Interference Rate (VIR) which measures the tendency of the model to hallucinate text from the slide that was never spoken.
+We evaluate four representative OLLMs: Qwen2.5-Omni (7B/3B) Xu et al.
+(2025a), MiniCPM-o-2.6 Yao et al.
+(2024), and Megrez-Omni Li et al.
+(2025a) on the SlideSpeech Wang et al.
+(2024b) dataset.
+The calculation follows a set-difference logic:
+  • Step 1: Context Isolation.
+Let $V_{slide}$ be the set of words on the slide and $V_{audio}$ be the set of words in the ground-truth transcript.
+We first identify the slide-exclusive vocabulary $V_{exclusive}=V_{slide}\setminus V_{audio}$.
+These are words that appear visually but are not spoken.
+  • Step 2: Interference Detection. Let $V_{pred}$ be the set of words in the model’s prediction. We calculate the intersection $I=V_{pred}\cap V_{exclusive}$.
+  • Step 3: Metric Calculation. If $I\neq\emptyset$, the sample is flagged as exhibiting visual interference. The VIR is defined as the percentage of such flagged samples in the dataset.
+The quantitative results are presented in Table 1.
+All evaluated OLLMs exhibit a high VIR, ranging from 12.87% to as high as 63.28% on the test set.
+Even the strongest baseline, Qwen2.5-Omni-7B, fails to suppress visual interference in over 12% of the samples.
+This consistent failure across different models indicates that this is a widespread issue rather than an isolated case.
+This phenomenon highlights a fundamental limitation in current paradigms: OLLMs lack an explicit mechanism to temporally decouple visual perception from auditory processing.
+Table 1: Comparison of Visual Interference Rate across different OLLMs on the SlideSpeech dataset.
+Model Dev set Test set
+Num=1,801 Num=3,053
+MiniCPM-o-2.6 57.96% 63.28%
+Megrez-Omni 45.14% 44.90%
+Qwen2.5-Omni-3B 15.43% 16.54%
+Qwen2.5-Omni-7B 13.71% 12.87%
+3 Method
+Figure 3: Overview of the Visually-Anchored Policy Optimization (VAPO) framework.
+The model takes audio, slide, and instruction as inputs and generates a structured <think><answer> sequence.
+In the <think> phase, the model extracts visual context, which serves as a semantic anchor (indicated by the red box) to guide the transcription in the <answer> phase.
+The policy is optimized via four rewards (Format, OCR, ASR, and Visual Anchoring).
+To mitigate the Visual Interference issue identified in Sec.
+2, we propose Visually-Anchored Policy Optimization (VAPO).
+Inspired by the human perception process, VAPO addresses this limitation by reshaping the model’s inference process to achieve a temporal decoupling of modalities.
+Specifically, it establishes a structured <think><answer> inference chain: the model first establishes a visual prior in the <think> phase, and then references this content as an anchor to guide speech transcription in the <answer> phase.
+This mechanism mimics the human-like “Look-then-Listen” workflow.
+The overall framework is illustrated in Fig.
+3.
+3.1 The “Look-then-Listen” Inference Chain
+We design a structured output format to explicitly enforce the necessary temporal decoupling between visual perception and auditory processing.
+Drawing inspiration from Chain-of-Thought (CoT) reasoning Jaech et al.
+(2024); Ma et al.
+(2025), the model is mandated to generate its output within a sequential <think><answer> structure.
+The <think> block corresponds to the “Look” phase: the model first processes the visual input and is tasked with extracting textual information from the slide image.
+This operation establishes a critical visual context prior before the auditory processing begins.
+Subsequently, the <answer> block corresponds to the “Listen” phase: the model generates the final transcription.
+In this phase, the model is required to reference the content anchored in the <think> block.
+This mechanism allows specialized terms from the slide to serve as semantic anchors, thereby assisting the model in resolving ambiguous audio and enhancing overall transcription accuracy.
+3.2 Multi-Objective Policy Optimization
+To make the model follow the proposed inference chain, we design four complementary reward functions to guide the model’s learning.
+We employ the Group Relative Policy Optimization (GRPO) algorithm Shao et al.
+(2024b) for model training.
+Format Reward.
+This reward aims to ensure that the model’s output strictly adheres to the <think><answer> format.
+A positive reward is assigned only if the model generates the complete structural tags.
+The reward function is defined as:
+$R_{\text{Format}}=\begin{cases}1,&\text{if the format is correct,}\\ 0,&\text{otherwise.}\end{cases}$ (1)
+OCR Reward.
+This reward evaluates the visual perception quality in the “Look” phase by comparing the text generated in the <think> block ($T_{t}$) with the ground truth slide text ($T_{s}$).
+We use the Word Error Rate (WER) as the metric, treating each Chinese character as a word where applicable.
+The reward is normalized and clipped to ensure non-negativity:
+$R_{\text{OCR}}=\max(1-\text{WER}(T_{t},T_{s}),0).$ (2)
+ASR Reward.
+This reward assesses the transcription quality in the “Listen” phase by comparing the output in the <answer> block ($T_{a}$) with the ground truth speech transcription ($T_{g}$).
+Similarly, it is derived from the WER:
+$R_{\text{ASR}}=\max(1-\text{WER}(T_{a},T_{g}),0).$ (3)
+Visual Anchoring Reward.
+This reward bridges the “Look” and “Listen” phases by incentivizing the model to ground spoken entities in the visual context.
+Let $E_{target}$ be the set of critical entities present in both the ground truth slide and the ground truth speech transcript.
+Let $E_{gen}$ be the subset of $E_{target}$ that successfully appears in both the generated <think> and <answer> blocks.
+The reward is defined as recall of these target anchors:
+$R_{\text{VA}}=\begin{cases}\frac{|E_{gen}|}{|E_{target}|},&\text{if }|E_{target}|>0,\\ 0,&\text{otherwise.}\end{cases}$ (4)
+This formulation encourages the model to explicitly capture and utilize the visual cues that are relevant to the current speech.
+Finally, the total reward is defined as a weighted sum of the four components, where $\lambda_{i}$ denotes the weighting hyperparameter for each reward:
+$\begin{split}R_{\text{total}}=\;&\lambda_{1}R_{\text{Format}}+\lambda_{2}R_{\text{OCR}}\\ &+\lambda_{3}R_{\text{ASR}}+\lambda_{4}R_{\text{VA}}.\end{split}$ (5)
+4 SlideASR-Bench: A Benchmark for Entity-rich SlideASR Task
+Table 2: Results on the SlideSpeech, a real-world English SlideASR dataset.
+$\dagger$ represents results from the original paper.
+The best and second-best results are in bold and underlined, respectively.
+Model Dev set Test set
+[…]
+Contextless
+Qwen2-Audio 12.56 12.85 8.72 91.43 13.19 13.59 7.53 92.91
+MiniCPM-o-2.6 16.09 16.68 8.14 91.98 18.71 19.41 8.90 91.50
+Qwen2.5-Omni-3B 15.53 16.22 6.30 93.76 12.00 12.45 5.72 94.41
+Qwen2.5-Omni-7B 11.75 12.20 5.39 94.78 11.75 12.20 5.39 94.78
+Qwen3-Omni-30B-A3B 10.87 11.31 5.02 95.04 11.71 12.21 4.64 95.50
+Slide text as context (Pipeline)
+Qwen2-Audio 139.81 145.05 69.94 85.40 146.08 152.41 56.99 88.98
+Mi-Dasheng 33.67 35.18 13.56 93.02 47.21 49.34 17.21 91.00
+Qwen3-Omni-30B-A3B 50.43 52.85 18.05 96.45 57.12 59.27 26.75 96.34
+[…]
+[…]
+Slide image as context (End-to-End)
+MiniCPM-o-2.6 182.96 192.83 51.07 86.26 210.37 220.96 60.92 83.22
+Qwen2.5-Omni-3B 12.22 12.74 5.26 95.17 19.99 20.71 9.80 94.44
+Qwen2.5-Omni-7B 13.65 14.13 7.19 92.84 14.97 15.58 6.33 93.99
+Qwen3-Omni-30B-A3B 19.85 20.64 9.30 95.59 24.13 24.88 13.44 94.74
+VAPO-3B (Ours) 9.84 10.31 3.61 96.54 10.73 11.24 3.55 96.57
+VAPO-7B (Ours) 8.62 9.08 2.48 97.61 10.31 10.84 2.87 97.32
+Our primary goal is to enhance ASR accuracy for domain-specific entities within visual presentation contexts.
+While existing datasets like SlideSpeech Wang et al.
 (2024b) and ChineseLips Zhao et al.
-(2025) lack a sufficient number of specialized named entities.
-To address this, we constructed SlideASR-Bench, a benchmark for the entity-rich SlideASR task, which comprises two subsets: SlideASR-S and SlideASR-R.
-SlideASR-S is built by synthesizing slides based on entity information from the ContextASR-Bench Wang et al.
+(2025), provide valuable benchmarks for general-domain scenarios, we observe that they often lack a sufficient density of domain-specific named entities.
+This scarcity creates a significant bottleneck for both training visually-anchored models and evaluating their capabilities in specialized scenarios.
+To bridge this gap, we introduce SlideASR-Bench, which comprises two distinct subsets: SlideASR-S, a large-scale synthetic corpus for training and evaluation, and SlideASR-R, a small real-world test set for stress-test.
+Detailed statistics are presented in Table 7.
+SlideASR-S.
+To train and evaluate models for entity-rich scenarios, we constructed SlideASR-S by extending the ContextASR-Bench Wang et al.
 (2025) dataset.
-To evaluate model performance in real-world scenarios, we additionally curated SlideASR-R by manually collecting 60 challenging samples from authentic professional reports across four domains: Chemistry, Medicine, Biology, and Artificial Intelligence.
-See Sec 5 for details.
-We conduct experiments on SlideSpeech Wang et al.
-(2024b), ChineseLips Zhao et al.
-(2025) and SlideASR-Bench.
-Results show that our proposed approach outperforms the state-of-the-art models, e.g., Qwen3-Omni-30B-A3B Xu et al.
-(2025b), particularly on entity-related metrics, while maintaining accuracy on non-entity text.
-Ablation studies show that each reward function in VAPO is essential for achieving optimal performanc.
-Attention visualization shows that the VAPO model follows the "Look before Transcription" procedure.
-The contributions can be summarized as follows:
-  • To the best of our knowledge, this work is the first to identify and analyze the limitations of OLLMs when applied to the SlideASR task.
-  • We introduce VAPO, a novel post-training method designed to improve the performance of OLLMs in the SlideASR task by enforcing a structured "Look before Transcription" reasoning process.
-  • We construct SlideASR-Bench, which consists of two dedicated datasets, SlideASR-S and SlideASR-R, to address the scarcity of domain-specific entities and provide robust benchmarks for the SlideASR task.
-The data and code will be released.
+ContextASR-Bench leverages LLMs, such as DeepSeek-R1 Guo et al.
+(2025), to generate colloquial text rich in named entities based on seed text.
+The seed text is sourced from Named Entity Recognition (NER) datasets across multiple domains (e.g., medicine, culture, and ecology).
+Text-to-speech models Du et al.
+(2024) are then used to convert the generated text into natural and fluent speech.
+We extract metadata for each sample from ContextASR-Bench, including the domain label $L_{\text{domain}}$ and the list of domain-specific entities $E$.
+Using the $L_{\text{domain}}$ and $E$ as input, we employ an LLM (e.g., Qwen2.5-14B-Instruct [^1] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct Team and others (2024)) to generate a short non-colloquial, formal text in slide style.
+The prompt guides the LLM to include a title and key points, ensuring that all entities from the original audio are naturally embedded in the generated text.
+Finally, we utilize Python’s Matplotlib library to render the generated text into slide images, generating a total of 8,467 samples (6,413 for training set, 2,054 for test set).
+We provide the prompt and data example in Appendix  A.
+SlideASR-R.
+Furthermore, to assess the model’s generalization ability in real, complex environments, we manually constructed a small-scale, high-quality, and challenging test set.
+We collected 60 real presentation audio clips and corresponding slide images from publicly available academic report videos, covering four specialized domains: chemistry, medicine, biology, and artificial intelligence.
+For each sample, we manually annotated the data by carefully comparing the speech and slide image, identifying the domain-specific entities that appear in both.
+We named this dataset SlideASR-R, which contains 200 domain-specific entities from real-world scenarios.
+5 Experiment
+5.1 Experimental Setup
+Table 3: Results on the SlideASR-Bench. The best and second-best results are in bold and underlined, respectively.
+Model SlideASR-S (en) SlideASR-S (zh) SlideASR-R
+[…]
+Contextless
+Qwen2-Audio 11.90 36.29 47.84 6.02 22.83 40.36 74.56 76.73
+MiniCPM-o-2.6 11.19 27.51 30.93 10.35 25.00 41.62 55.85 65.37
+Qwen2.5-Omni-3B 8.37 24.15 31.04 4.47 19.89 38.08 61.31 66.83
+Qwen2.5-Omni-7B 8.15 23.44 27.77 4.34 17.54 32.80 53.68 63.37
+Qwen3-Omni-30B-A3B 9.06 14.61 15.53 20.77 23.31 22.49 40.43 41.09
+Slide text as context (Pipeline)
+Qwen2-Audio 92.16 66.38 24.82 39.09 50.58 31.52 59.04 21.29
+Mi-Dasheng 78.98 49.85 30.58 66.88 56.30 32.30 47.52 26.73
+Qwen3-Omni-30B-A3B 34.65 32.35 8.56 9.76 15.85 13.54 34.01 28.22
+Slide image as context (End-to-End)
+MiniCPM-o-2.6 112.90 49.65 15.01 89.53 61.25 45.67 63.73 66.83
+Qwen2.5-Omni-3B 100.08 53.19 18.72 86.86 65.62 9.62 49.00 53.47
+Qwen2.5-Omni-7B 57.21 35.76 15.04 91.83 54.04 3.36 41.77 35.15
+Qwen3-Omni-30B-A3B 101.45 59.64 12.08 79.21 46.45 5.54 32.26 24.75
+VAPO-3B (Ours) 4.90 3.19 3.73 2.47 4.21 2.22 27.28 19.31
+VAPO-7B (Ours) 4.60 2.83 2.97 2.13 3.78 1.36 26.48 15.35
+[…]
+Baselines & Settings.
+We benchmark mainstream LALMs: Qwen2-Audio Chu et al.
+(2024) and Mi-Dasheng Dinkel et al.
+(2025) and OLLMs: MiniCPM-o-2.6 Yao et al.
+(2024) Qwen2.5-Omni Xu et al.
+(2025a), Qwen3-Omni Xu et al.
+(2025b) on SlideSpeech Wang et al.
+(2024b) and SlideASR-Bench across three settings:
+  1. 1.
+    Contextless, which only use audio as inputs;
+  2. 2.
+    Slide text as context (Pipeline), employing PaddleOCR Cui et al. (2025) for textual context extraction following Zhao et al. (2025);
+  3. 3.
+    Slide image as context (End-to-end).
+We provide more evaluation details in Appendix B. Additionally, we report results on the ChineseLips Zhao et al. (2025) dataset in Appendix C to demonstrate the generalization of VAPO.
+Metrics.
+We utilize two sets of metrics consistent with prior works.
+For SlideSpeech Wang et al.
+(2024b), we employ four metrics: WER, which measures overall transcription; U-WER, the unbiased WER on non-keyword segments for general transcription quality; B-WER, the biased WER focusing on keyword spans; and Recall, which is the percentage of correctly recognized keywords.
+Furthermore, for SlideASR-Bench, consistent with ContextASR-Bench Wang et al.
+(2025), we focus on three metrics: WER; NE-WER, the WER of the named entity portion; and NE-FNR, the False Negative Ratio of named entities, which measures the proportion of missed ground-truth entities.
+See Appendix B.2 for specific calculation details of the metrics.
+5.2 Main Results and Analysis
+Results on SlideSpeech.
+Table 2 presents the results on the real-world SlideSpeech dataset.
+Ideally, visual context should enhance recognition; however, we observe a detrimental effect in baseline models.
+Most baselines exhibit performance degradation when incorporating slide information compared to the contextless setting.
+For instance, the Qwen3-Omni Xu et al.
+(2025b) shows increased WER in both pipeline and end-to-end settings compared to using audio alone.
+In stark contrast, our VAPO method effectively leverages visual cues, achieving the best performance.
+VAPO-7B reaches a WER of 10.31 and a Recall of 97.32, significantly outperforming the Qwen3-Omni baseline and the previous SOTA MaLa-ASR Yang et al.
+(2024).
+Results on SlideASR-Bench. Table 3 reports the performance on our benchmark. Two critical observations highlight the superiority of VAPO:
+1) Audio-only models struggle with specialized entities.
+On the challenging SlideASR-R subset, even the strongest audio-only model (Qwen3-Omni) suffers from a high NE-FNR of 41.09, underscoring the necessity of visual presentation context for domain-specific terms.
+2) Naive visual integration leads to catastrophic failure.
+On SlideASR-S, baseline OLLMs fail to utilize visual context effectively.
+Instead of improving, they suffer from severe visual interference.
+For example, Qwen3-Omni in the end-to-end setting yields an exploded WER of 101.45 and a worsened NE-WER of 59.64 (vs.
+14.61 in audio-only), indicating that the model is hallucinating slide content rather than transcribing speech.
+Conversely, on SlideASR-S (en), VAPO-3B achieves a remarkable WER of 4.90.
+Furthermore, on the SlideASR-R, VAPO-7B reduces the NE-FNR from the best baseline’s 28.22 to 15.35.
+These results confirm that VAPO successfully enforces the “Look-then-Listen” paradigm, accurately anchoring visual entities without succumbing to interference.
+We present case study in Appendix D to provide visual comparison.
+Table 4: Ablation results of rewards on SlideASR-R.
+[…]
+Qwen2.5-Omni-3B
+49.00 53.47
+37.23 31.19
+29.97 22.28
+27.28 19.31
+Qwen2.5-Omni-7B
+41.77 35.15
+28.63 20.30
+26.75 18.32
+26.48 15.35
+Table 5: Sensitivity analysis of reward weights ($\lambda_{1}$$\lambda_{2}$$\lambda_{3}$$\lambda_{4}$) on SlideASR-S.
+The balanced 1:1:1:1 strategy achieves the best robustness across metrics (N-W: NE-WER, N-F: NE-FNR).
+Weights SlideASR-S (en) SlideASR-S (zh)
+WER N-W N-F WER N-W N-F
+1:1:1:1 4.90 3.19 3.73 2.47 4.21 2.22
+1:1:1:2 5.27 3.34 3.78 2.50 4.30 2.09
+1:1:2:1 5.32 4.12 3.91 2.48 4.38 2.09
+1:2:1:1 5.17 3.45 3.80 2.51 4.23 1.99
+5.3 Ablation Study
+Ablation on Reward Functions.
+Table 4 details the stepwise contributions on SlideASR-R.
+Compared to the baseline, introducing the ASR Reward yields the most significant boost, drastically reducing NE-WER by stabilizing generation.
+Adding the OCR Reward further refines the visual prior quality.
+Finally, incorporating the Visual Anchoring Reward achieves the best performance.
+This confirms that while ASR and OCR guarantees individual modality perception, the VA reward is essential for effective referencing.
+We provide additional ablation results in Appendix E.
+Sensitivity to Reward Weights.
+To investigate hyperparameter sensitivity, we evaluate varying weight configurations on SlideASR-S, as detailed in Table 5.
+The results confirm that the balanced 1:1:1:1 scheme yields the most robust overall performance.
+We observe an inherent trade-off: doubling the VA reward ($\lambda_{4}$=2) enhances entity recall on the Chinese subset (improving NE-FNR to 2.09), but at the cost of increased overall WER due to over-aggressive anchoring.
+Conversely, double the ASR reward ($\lambda_{3}$=2) proves counterproductive, suppressing the model’s reliance on visual cues and significantly degrading NE-WER (e.g., rising from 3.19 to 4.12 on the English subset).
+Consequently, we adopt the equal weighting strategy for its simplicity and effectiveness.
+5.4 Robustness against Mismatched Slides
+To verify that VAPO performs valid visual anchoring rather than blindly copying slide content, we evaluated the model’s performance when fed with randomly mismatched slide images.
+As illustrated in Fig.
+4, the model exhibits strong robustness: the overall WER remains stable at 6.70%, which is comparable to the audio-only baselines, including Qwen2.5-Omni-3B (6.42%) and the audio-only variant of VAPO-3B (7.40%).
+This indicates that VAPO effectively mitigates visual interference, safely falling back to auditory perception when visual context is unreliable.
+Figure 4: Robustness analysis of VAPO-3B under mismatched visual context on SlideASR-S.
+Table 6: Comparison between VAPO and SFT baselines on SlideASR-S.
+"SFT w/o think" is trained to directly predict the transcription, whereas "SFT w/ think" is supervised using the structured inference chain.
+[…]
+Qwen2.5-Omni-3B
++ SFT w/o think 8.44 8.18 9.30
++ SFT w/ think 6.60 6.47 7.10
++ VAPO 3.88 3.84 3.00
+Qwen2.5-Omni-7B
++ SFT w/o think 10.58 10.51 12.77
++ SFT w/ think 6.73 5.56 6.53
++ VAPO 3.37 3.07 2.00
+Figure 5: Attention visualization of VAPO-7B.
+The dense orange connections demonstrate the visual anchoring mechanism, showing that the model explicitly attends to extracted entities in <think> to guide the final transcription.
+5.5 Impact of Training Paradigm: VAPO vs. Supervised Fine-Tuning (SFT)
+To disentangle the contributions of the structured format from the reinforcement learning (RL) optimization, we compare three training settings on SlideASR-S: SFT w/o think (direct mapping), SFT w/ think (structured but SFT-only), and VAPO.
+Results in Table 6 highlights two key findings:
+1) Efficacy of the Inference Chain.
+Introducing a structured inference chain yields significant gains, reducing the WER of Qwen2.5-Omni-7B from 10.58 to 6.73.
+This confirms that temporally decoupling the “Look” and “Listen” phases inherently mitigates visual interference.
+2) Necessity of RL.
+VAPO further lowers the WER to 3.37.
+This indicates that while SFT teaches the format, it fails to establish a sufficient connection between <think> and <answer>.
+VAPO, driven by the multi-dimensional rewards, ensures the <think> block is effectively utilized for the final transcription.
+5.6 Attention Visualization
+To empirically validate the effectiveness of the visual anchoring mechanism, we visualize the attention weights of VAPO-7B on a sample from SlideSpeech, as shown in Fig.
+5.
+The visualization reveals a distinct semantic reference pattern: during the generation of the <answer> block, the model heavily attends to the corresponding extracted text within the <think> block.
+Specifically, take the proper noun “Concentriq” as an example.
+After generating the token “Concent” in <answer>, the model pays significant attention to the “ri” token in <think> and subsequently generates it.
+It then refers to the “q” token in the <think> block, enabling accurate and complete transcription of “Concentriq”.
+A similar process occurs for the entity “proscia”.
+This qualitative evidence confirms that VAPO successfully enforces the “Look-then-Listen” paradigm, where the model explicitly references the visual context prior.
+We provide more cases of attention visualization in Appendix F.
 6 Related Works
 Contextual ASR.
-The objective of contextual ASR is to incorporate contextual information, including domain labels, entity lists, and conversational history, into the speech recognition system in order to improve the recognition accuracy of named entities, and domain-specific terminology Bai et al.
+The objective of contextual ASR is to incorporate contextual information, including domain labels, entity lists, and conversational history, into the speech recognition system in order to improve the recognition accuracy of domain-specific terminology Bai et al.
 (2024); Xiao et al.
@@
 Omni-modal Large Language Models.
-Recently, OLLMs Fu et al.
+Recently, OLLMs Hurst et al.
+(2024); Fu et al.
 (2025); Yao et al.
@@
 Benefiting from the unified modeling capability across visual and audio modalities, they are expected to solve the SlideASR task in an end-to-end manner.
-However, in practical scenarios, the models exhibit unstable behavior, for instance, they sometimes reproduce the textual content from the slides instead of generating the expected speech transcription.
+However, in practical scenarios, the models exhibit visual interference, for instance, they sometimes reproduce the textual content from the slides instead of generating the expected speech transcription.
 Chain-of-Thought Reasoning.
-CoT reasoning is a breakthrough approach that enhances the complex reasoning capabilities of LLMs.
+CoT reasoning is a breakthrough approach that enhances the reasoning capabilities of LLMs.
 Recent works Jaech et al.
@@
 (2024); Ma et al.
+(2025); Lin et al.
 (2025); Diao et al.
 (2025), demonstrating the general effectiveness of making reasoning processes explicit.
-3 The Failure of OLLMs in SlideASR
-3.1 Problem Formulation
-Figure 2: Comparison of OLLM outputs with and without slide context.
-We formally define the SlideASR task as follows.
-Given a speech signal $A$ and a slide image $I$ synchronized with the speech, our goal is to train a model $f_{\theta}$ that takes $A$ and $I$ as joint inputs and generates the most likely corresponding text transcription $Y=\{y_{1},y_{2},…,y_{n}\}$, where $y_{n}$ is the $n$-th token in the text sequence.
-The objective of this task can be expressed as maximizing the conditional probability $P(Y|A,I;\theta)$, where $\theta$ represents the learnable parameters of the model.
-$\theta^{*}=\arg\max_{\theta}P(Y|A,I;\theta)$ (1)
-In our work, $f_{\theta}$ is an OLLM is that takes both image and audio as input and generates the final text $Y$ in an auto-regressive manner.
-3.2 Observations of OLLMs on SlideASR
-Although OLLMs can process both auditory and visual modalities simultaneously, we observe that their behavior on the SlideASR task is unstable.
-Fig 2 shows an example.
-We selected a sample from SlideSpeech Wang et al.
-(2024b) and had Qwen2.5-Omni-7B transcribe the speech under two conditions: audio-only and audio with slide image.
-When the model receives only audio, it produces correct ASR results.
-However, when both audio and the corresponding slide image are provided, the model ignores the audio signal and instead outputs the text content from the slide image, behaving like a OCR system.
-To quantitatively demonstrate the instability of OLLMs’ behavior, we calculated the proportion of OCR-like behavior exhibited by four models Qwen2.5-Omni-7B Xu et al.
-(2025a), Qwen2.5-Omni-3B Xu et al.
-(2025a), MiniCPM-o-2.6 Yao et al.
-(2024), and Megrez-Omni Li et al.
-(2025a) on the SlideSpeech dataset.
-The judgment process involves three steps.
-1) Identify Common Vocabulary: We first identify the words common to both the ground truth speech transcription and the slide text, denoted as $V_{\text{common}}$.
-2) Isolate Slide-Only Vocabulary: We then create a set of slide-only words, $V_{\text{slide\_only}}$, by removing the common words from the slide’s full vocabulary.
-These words are only accessible to the model via OCR.
-3) Detect OCR Behavior: Finally, we check for any intersection between the model’s output and $V_{\text{slide\_only}}$.
-If such an intersection exists, it indicates that the model has exhibited OCR behavior.
-Table 1: Percentage of samples with OCR behavior on SlideSpeech dataset.
-Dev set Test set
-Num=1,801 Num=3,053
-MiniCPM-o-2.6 57.96% 63.28%
-Megrez-Omni 45.14% 44.90%
-Qwen2.5-Omni-3B 15.43% 16.54%
-Qwen2.5-Omni-7B 13.71% 12.87%
-The results are shown in Table 1.
-All models exhibited a significant proportion of OCR behavior on the dataset, generating words present on the slide but absent in the audio.
-This fundamental and widespread failure demonstrates that simply prompting OLLMs is insufficient and that a new, more structured approach is required to control their reasoning process.
-4 Visually-Anchored Policy Optimization
-To address the issue of uncontrollable behavior in OLLMs for the SlideASR task, we propose Visually-Anchored Policy Optimization (VAPO).
-The core idea of VAPO is to transform the model from an unreliable "black-box" into a controllable one that follows a structured and effective reasoning path.
-This is achieved through explicit structured reasoning and multi-objective reward-based policy optimization.
-The process of VAPO is shown in Fig 3.
-Figure 3: Overview of the Visually-Anchored Policy Optimization framework.
-The OLLM takes audio, slide, and instruction as input, generates a structured output, and is optimized via reward functions that guide the policy update.
-4.1 Structured Reasoning Format
-Inspired by CoT Reasoning Jaech et al.
-(2024), we design a mandatory output format to guide the model in following the reasoning principle we call "Look before Transcription".
-Specifically, the model is required to generate its output within a unified structure of <think><answer>.
-In the <think> block, the model first processes and digests the visual information.
-Before generating the final transcription, it must output recognized text from the slide image $I$.
-This step acts as an explicit internal OCR, ensuring the model sees and understands key slide information before processing the audio.
-The <answer> block is where the final transcription $Y$ is generated.
-Instead of generating directly, the model is required to reference the visual content anchored in the <think> block.
-This allows specialized terms and technical jargon from the slide to serve as prior knowledge, helping the model resolve ambiguous audio and homophones, improving the transcription of key named entities.
-Through this format, we decompose the complex and implicit SlideASR task into an ordered, two-stage explicit reasoning process.
-4.2 Reinforcement Learning Optimization
-We design four reward functions to guide the model’s learning comprehensively from different dimensions.
-To optimize the model’s policy with these rewards, we employ the Generative Representational Policy Optimization (GRPO) Shao et al.
-(2024b) algorithm for fine-tuning the model.
-Format Reward.
-This reward aims to ensure that the model’s output strictly adheres to the <think></think><answer></answer> format.
-A positive reward is given if the model generates the complete structure.
-The reward function is as follows:
-$R_{\text{Format}}=\begin{cases}1,&\text{If the format is correct}\\ 0,&\text{otherwise.}\end{cases}$ (2)
-OCR Reward.
-This reward is used to evaluate the accuracy of the content in the <think> block.
-We compare the text generated by the model in <think> with the actual text on the slide.
-We calculate the reward using the Word Error Rate (WER).
-If the sample is in Chinese, each character is treated as a word.
-To ensure the reward is non-negative, we apply a clipping mechanism that prevents it from going below zero.
-We denote the text in the <think> block as $T_{t}$, and the text on the slide as $T_{s}$.
-The reward function is as follows:
-$R_{\text{OCR}}=max(1-\text{WER}(T_{t},T_{s}),0)$ (3)
-ASR Reward.
-This reward is used to evaluate the overall quality of the final transcription text in the <answer> block.
-We compare the output in <answer> with the ground truth speech transcription.
-Similarly, we apply the same clipping mechanism in ASR reward.
-We denote the text in the <answer> as $T_{a}$ block, and the ground truth speech transcription as $T_{g}$.
-The reward function is as follows:
-$R_{\text{ASR}}=max(1-\text{WER}(T_{a},T_{g}),0)$ (4)
-Visual Anchoring Reward.
-This reward in the VAPO is responsible for establishing a connection between the <think> and <answer> blocks.
-This reward specifically incentivizes the model to correctly use the named entities recognized in the <think> block within the <answer> block.
-The calculation process is as follows.
-First, we extract the set of correctly identified entities $E_{\text{think}}$ from the output of the <think>.
-Then, we calculate the F1 score of this entity set $E_{\text{think}}$ in the output of the <answer>.
-The reward function is as follows:
-$R_{\text{VA}}=F1_{score}(E_{think},T_{answer}).$ (5)
-This reward encourages the model to reference and utilize the slide information, effectively mitigating the issue of disjoint processing between the "seeing" and "hearing".
-Finally, the model’s total reward is defined as a weighted sum of the four rewards described above:
-$\begin{split}R_{total}=&\lambda_{1}R_{\text{Format}}+\lambda_{2}R_{\text{OCR}}\\ &+\lambda_{3}R_{\text{ASR}}+\lambda_{4}R_{\text{VA}}\end{split}$ (6)
-$\lambda$ denotes the weighting hyperparameters for each individual reward.
-5 SlideASR-Bench: A Benchmark for entity-rich SlideASR Task
-Our goal is to improve the accuracy of model transcription for domain-specific entities in the context of slides.
-Existing public SlideASR datasets, such as SlideSpeech Wang et al.
-(2024b) and ChineseLips Zhao et al.
-(2025), provide a valuable benchmark for real-world general-domain scenarios.
-However, we observe that these datasets often lack a sufficient density of domain-specific named entities in both speech and slides.
-This creates a significant challenge in training a model capable of handling entity-rich slides and speech, as well as objectively evaluating its performance in entity-rich SlideASR scenarios.
-To address this critical resource bottleneck, we constructed SlideASR-Bench, which includes two datasets: a synthetic dataset, SlideASR-S, for training and evaluation, and a challenging evaluation set, SlideASR-R, for assessing real-world performance.
-Table 2 presents the detailed information.
-SlideASR-S.
-To train and evaluate models for entity-rich scenarios, we constructed SlideASR-S based on the ContextASR-Bench Wang et al.
-(2025) dataset.
-ContextASR-Bench leverages LLMs, such as DeepSeek-R1 Guo et al.
-(2025), to generate colloquial text rich in named entities based on seed text.
-The seed text is sourced from Named Entity Recognition (NER) datasets across multiple domains, such as medicine, culture, ecology.
-Then, models such as CosyVoice2 Du et al.
-(2024) and XTTS-v2 Casanova et al.
-(2024), which are text-to-speech models, are used to convert the generated text into natural and fluent speech.
-We extract metadata for each sample from ContextASR-Bench, including the domain label $L_{\text{domain}}$ and the list of domain-specific entities $E=\{e_{1},e_{2},\dots,e_{k}\}$.
-Using the $L_{\text{domain}}$ and $E$ as input, we employ a LLM (e.g., Qwen2.5-14B-Instruct  [^1] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct) to generate a short descriptive text in slide style.
-The prompt guides the LLM to include a title and key points, ensuring that all entities from the original audio are naturally embedded in the generated text.
-The prompt can be found in Appendix  A.
-Finally, we use Python’s Matplotlib library to render the text into an image, simulating a presentation slide.
-Each generated image serves as synthetic visual context.
-SlideASR-R.
-Although synthetic data is useful for model training and evaluation, there remains a certain domain gap between synthetic and real-world scenarios.
-To assess the model’s generalization ability in real, complex environments, we manually constructed a small-scale, high-quality, and challenging test set.
-We collected 60 real presentation audio clips and corresponding slide images from publicly available academic report videos, covering four specialized domains: chemistry, medicine, biology, and artificial intelligence.
-For each sample, we manually annotated the data by carefully comparing the speech and slide image, identifying the domain-specific entities that appear in both.
-We named this dataset SlideASR-R (R for Real), which contains 200 domain-specific entities from real-world scenarios.
-Despite its relatively small size, it provides a highly challenging benchmark for assessing model performance in practical applications.
-Table 2: Details of our proposed SlideASR-Bench.
-#Sample #Entity #Hour
-SlideASR-S (Train set) 6,413 44,240 67.3
-SlideASR-S (Test set) 2,054 13,895 18.5
-SlideASR-R 60 200 0.35
-6 Experiment
-6.1 Setup
-Table 3: Results on the SlideSpeech, a real-world English SlideASR dataset.
-$\dagger$ represents results from the original paper.
-The best and second-best results are in bold and underlined, respectively.
-Model Dev Test
-[…]
-Contextless
-Qwen2-Audio 12.56 12.85 8.72 91.43 13.19 13.59 7.53 92.91
-Mi-Dasheng 13.63 13.97 9.06 90.97 14.61 15.04 8.52 91.59
-MiniCPM-o-2.6 16.09 16.68 8.14 91.98 18.71 19.41 8.90 91.50
-Qwen2.5-Omni-3B 15.53 16.22 6.30 93.76 12.00 12.45 5.72 94.41
-Qwen2.5-Omni-7B 11.75 12.20 5.39 94.78 11.75 12.20 5.39 94.78
-Qwen3-Omni-30B-A3B 10.87 11.31 5.02 95.04 11.71 12.21 4.64 95.50
-Slide text as context (Pipeline)
-Qwen2-Audio 139.81 145.05 69.94 85.40 146.08 152.41 56.99 88.98
-Mi-Dasheng 33.67 35.18 13.56 93.02 47.21 49.34 17.21 91.00
-Qwen3-Omni-30B-A3B 50.43 52.85 18.05 96.45 57.12 59.27 26.75 96.34
-[…]
-[…]
-Slide image as context (End-to-End)
-MiniCPM-o-2.6 182.96 192.83 51.07 86.26 210.37 220.96 60.92 83.22
-Qwen2.5-Omni-3B 12.22 12.74 5.26 95.17 19.99 20.71 9.80 94.44
-Qwen2.5-Omni-7B 13.65 14.13 7.19 92.84 14.97 15.58 6.33 93.99
-Qwen3-Omni-30B-A3B 19.85 20.64 9.30 95.59 24.13 24.88 13.44 94.74
-VAPO-3B (Ours) 9.84 10.31 3.61 96.54 10.73 11.24 3.55 96.57
-VAPO-7B (Ours) 8.62 9.08 2.48 97.61 10.31 10.84 2.87 97.32
-Table 4: Results on the SlideASR-Bench. The best and second-best results are in bold and underlined, respectively.
-Model SlideASR-S (en) SlideASR-S (zh) SlideASR-R
-WER NE-WER NE-FNR WER NE-WER NE-FNR NE-WER NE-FNR
-Contextless
-Qwen2-Audio 11.90 36.29 47.84 6.02 22.83 40.36 74.56 76.73
-Mi-Dasheng 12.16 30.32 35.40 4.67 19.33 35.81 54.08 61.39
-MiniCPM-o-2.6 11.19 27.51 30.93 10.35 25.00 41.62 55.85 65.37
-Qwen2.5-Omni-3B 8.37 24.15 31.04 4.47 19.89 38.08 61.31 66.83
-Qwen2.5-Omni-7B 8.15 23.44 27.77 4.34 17.54 32.80 53.68 63.37
-Qwen3-Omni-30B-A3B 9.06 14.61 15.53 20.77 23.31 22.49 40.43 41.09
-Slide text as context (Pipeline)
-Qwen2-Audio 92.16 66.38 24.82 39.09 50.58 31.52 59.04 21.29
-Mi-Dasheng 78.98 49.85 30.58 66.88 56.30 32.30 47.52 26.73
-Qwen3-Omni-30B-A3B 34.65 32.35 8.56 9.76 15.85 13.54 34.01 28.22
-Slide image as context (End-to-End)
-MiniCPM-o-2.6 112.90 49.65 15.01 89.53 61.25 45.67 63.73 66.83
-Qwen2.5-Omni-3B 100.08 53.19 18.72 86.86 65.62 9.62 49.00 53.47
-Qwen2.5-Omni-7B 57.21 35.76 15.04 91.83 54.04 3.36 41.77 35.15
-Qwen3-Omni-30B-A3B 101.45 59.64 12.08 79.21 46.45 5.54 32.26 24.75
-VAPO-3B (Ours) 4.90 3.19 3.73 2.47 4.21 2.22 27.28 19.31
-VAPO-7B (Ours) 4.60 2.83 2.97 2.13 3.78 1.36 26.48 15.35
-Figure 4: Attention visualization. Left: input image and transcribed audio text. Right: attention flows.
-We fine-tune the Qwen2.5-Omni-3B and Qwen2.5-Omni-7B models on the SlideASR-S training set using the proposed VAPO algorithm.
-All experiments were run on 4 NVIDIA A100 GPUs with 80 GB of memory each.
-The weights of the reward functions, $\lambda_{1}$ to $\lambda_{4}$, are all set to 1.
-We evaluate models on SlideSpeech Wang et al.
-(2024b) and SlideASR-Bench, using three settings: Contextless, where only audio is used; Slide text as context, a pipeline setting where OCR extracts text from slides and combines it with audio; and Slide image as context, an end-to-end setting that directly inputs both slide images and audio.
-We select mainstream LALMs Chu et al.
-(2024); Dinkel et al.
-(2025) and OLLMs Yao et al.
-(2024); Xu et al.
-(2025a, b) as baselines.
-Note that the same model may support multiple settings.
-For example, Qwen3-Omni-30B-A3B Xu et al.
-(2025b) can take audio alone as input or accept text or image as context.
-Following Zhao et al.
-(2025), we use PaddleOCR  [^2] https://github.com/PaddlePaddle/PaddleOCR.
-for OCR.
-See Appendix B for details of evaluation prompts, metrics and baseline models.
-Additionally, we present the results on ChineseLips Zhao et al.
-(2025), a real-world Chinese dataset, in Appendix C (Table 6) to demonstrate VAPO’s generalization in real-world scenarios.
-6.2 Main Results
-Results on SlideSpeech.
-Table 3 shows the results on the real-world SlideSpeech dataset.
-It can be observed that baseline models generally perform poorly when incorporating slide information, with performance even degrading compared to the contextless setting.
-For instance, the Qwen3-Omni-30B-A3B Xu et al.
-(2025b), using either slide text or the slide image as context exhibits a higher WER compared to using audio alone.
-In contrast, our proposed VAPO method achieves the best results on the SlideSpeech dataset.
-The VAPO-7B model reaches a WER of 10.31 and a recall of 97.32 on the test set, outperforming baselines such as Qwen3-Omni-30B-A3B (Recall=95.5) and MaLa-ASR Yang et al.
-(2024) (WER=11.26).
-Results on SlideASR-Bench.
-Table 4 presents the results on SlideASR-Bench.
-Unlike SlideSpeech, on SlideASR-Bench, all audio-only models struggle to recognize specialized named entities.
-For example, on the SlideASR-R dataset, even the strongest audio-only model, Qwen3-Omni-30B-A3B, reaches a NE-FNR of 41.09.
-This underscores the essential role of incorporating visual information in entity-rich ASR tasks.
-Additionally, baseline models fail to effectively utilize contextual information, whether in text or image form.
-For example, on SlideASR-S, when Qwen3-Omni-30B-A3B uses OCR text or slide images as context, although NE-WER decreases compared to audio-only input, the overall WER increases significantly.
-This aligns with our observations in Section 3, the model is not transcribing the speech properly but rather outputting the context content instead.
-The VAPO method achieves significant improvements in recognizing key entities.
-For example, on the challenging real-world evaluation set SlideASR-R, VAPO-7B reduces the NE-FNR from the baseline best of 28.22 to 15.35.
-These results indicate that VAPO can accurately extract and anchor entity text from slides, thereby improving transcription accuracy.
-See Appendix D, for the successful cases and analysis of a failure case.
-6.3 Ablation Results
-Table 5: Ablation results on the SlideASR-R dataset.
-ASR
-Reward
-OCR
-Reward
-VA
-Reward
-[…]
-Qwen2.5-Omni-3B
-49.00 53.47
-37.23 31.19
-29.97 22.28
-27.28 19.31
-Qwen2.5-Omni-7B
-41.77 35.15
-28.63 20.30
-26.75 18.32
-26.48 15.35
-Table 5 presents the ablation results for different reward functions.
-The results show that, the ASR reward enables the models to initially reference slide information.
-The OCR reward further improves the accuracy of slide text extraction, enhancing overall performance.
-Finally, adding the Visual Anchoring (VA) reward strengthens the models’ focus on key entities in <think>, achieving the best results.
-Additional ablation results for SlideSpeech (Table 7) and SlideASR-S (Table 8) are in Appendix E.1.
-Table 9 in Appendix E.2 presents the ablation results for the reward weights ($\lambda_{1}$$\lambda_{2}$$\lambda_{3}$$\lambda_{4}$), demonstrating the trade-off between overall transcription accuracy and entity recognition performance, and how different weight configurations impact this balance.
-The 1:1:1:1 configuration optimally balances both.
-6.4 Attention Visualization
-To examine the model’s behavior during final transcription, we visualize the attention weights of VAPO-7B on a SlideSpeech case, as shown in Fig 4.
-We observed that when the model generates speech transcription in <answer>, it refers to the entity information in <think>.
-For instance, for the key entity "Concentriq", after generating the token "Concent" <answer>, the model pays significant attention to the "ri" token in <think> and subsequently generates it.
-It then refers to the "q" token in the <think> block, enabling accurate and complete transcription of "Concentriq".
-A similar process occurs for the entity "proscia".
-This demonstrates that the model indeed references the slide information during transcription.
-Further cases can be found in Appendix F.
 7 Conclusion
-This paper identifies key failures in OLLMs when applied to the SlideASR task, where they often disregard audio input and function merely as OCR systems.
-To address this, we introduce Visually-Anchored Policy Optimization (VAPO), a novel training method that enforces a structured "Look before Transcription" reasoning process.
-By leveraging a <think><answer> format and multi-faceted reward functions, VAPO significantly enhances the model’s ability to integrate visual and auditory information.
-Furthermore, we developed SlideASR-Bench, an entity-rich benchmark, to facilitate more robust training and evaluation.
-Extensive experiments demonstrate that VAPO substantially improves the transcription accuracy, particularly for specialized terms, establishing a more effective end-to-end paradigm for SlideASR.
+In this work, we uncover a critical Visual Interference phenomenon in OLLMs applied to the SlideASR task, where models succumb to modality dominance and ignore auditory inputs.
+To address this, we introduce Visually-Anchored Policy Optimization (VAPO), a novel post-training method that enforces a human-like “Look-then-Listen” inference chain.
+By leveraging a structured <think><answer> format and multi-dimensional reward optimization, VAPO effectively decouples visual perception from auditory transcription.
+Furthermore, we contribute SlideASR-Bench, a comprehensive benchmark designed to address the data scarcity in entity-rich scenarios.
+Extensive experiments demonstrate that VAPO not only eliminates visual interference but also sets a new state-of-the-art performance, particularly in recognizing specialized domain entities.
 8 Limitations
@@
 Our current approach is highly specialized for leveraging textual information from presentation slides and does not incorporate other visual cues, such as images of entities (e.g., pictures of specific drugs).
-In the future, we will adapt our "Look before Transcription" paradigm to handle more diverse multimodal environments, where various visual elements play a crucial role.
+In the future, we will adapt our “Look-then-Listen” paradigm to handle more diverse multimodal environments, where various visual elements play a crucial role.
 Real-World Robustness.
@@
 Nonetheless, a subtle domain gap may still exist, as synthetic slides may not fully capture the stylistic diversity and visual noise (e.g., complex diagrams, low-quality images) of all real-world presentations.
-Inference Efficiency: The structured reasoning process of VAPO introduces a computational overhead, resulting in higher inference latency compared to models of the same size (detailed in Appendix G).
+Inference Efficiency.
+The structured reasoning process of VAPO introduces a computational overhead, resulting in higher inference latency compared to models of the same size (detailed in Appendix G).
 This makes VAPO most suitable for offline applications where accuracy is critical.
@@
 We advocate for responsible development and believe that human oversight is essential for any critical applications.
+10 Acknowledgments
+This work is supported by the National Key R&D Program of China (No.
+2023YFC3310700), the Open Funding Programs of State Key Laboratory of AI Safety, the National Natural Science Foundation of China (No.
+2576030), and the Beijing Nova Program under Grant 20250484899.
 Appendix
-Appendix A Prompts for Generate SlideASR-S
-The prompt for the LLM to generate text paragraphs based on a domain label and an entity list is as follows.
+Appendix A Details of SlideASR-Bench
+Statistics of SlideASR-Bench. Table 7 presents the key statistical metrics for SlideASR-Bench, including the number of samples, entities, and hours.
+Table 7: Details of our proposed SlideASR-Bench.
+Subset Sample Entity Hour
+SlideASR-S (Training set) 6,413 44,240 67.3
+SlideASR-S (Test set) 2,054 13,895 18.5
+SlideASR-R 60 200 0.35
+Prompt for Generating SlideASR-S. The prompt for the LLM to generate text paragraphs based on a domain label and an entity list is as follows. Fig. 6 shows an example of SlideASR-S.
 Prompt for Qwen2-14B-Instruct to generate slide text Given a domain label and a list of entities, generate a title and a paragraph for use in a PPT report, with the requirement that the paragraph includes these entities, Keep paragraphs within 150 words.
-Domain label:
-List of entities:
-Output format:
-Title
-Paragraph
+Domain label: {} List of entities: {} Output format: ### Title ### Paragraph
+Figure 6: Data example of SlideASR-S.
 Appendix B Evaluation Details
@@
 The speech is the speaker’s talk accompanied by a slide, with the text of the slide being: {}
-Transcribe the speech into text by integrating the speech with the slide content.
-### Slide image as context
+Transcribe the speech into text by integrating the speech with the slide content. ### Slide image as context
 Taking the image content into account, convert the audio to text.
@@
 Role:User
-First, recognize the text in the image and output it within <think> </think>. Then, referring to the thinking content, output the speech recognition result within <answer> </answer>
+First, recognize the text in the image and output it within <think> </think>. Then, referring to the thinking content, output the speech recognition result within <answer> </answer>.
 B.2 Metrics
@@
   • NE-WER: WER of named entity portion, we first perform a fuzzy match to identify key entities (with an edit distance tolerance of $\frac{2}{WordCountOfEntity}-1$) in the model’s output, and then calculate the WER based on the fuzzy-matched entities.
-  • NE-FNR: The false negative ratio of named entities, calculated as $1-\frac{N}{H}$, where $H$ and $N$ denote the recognized and ground-truth entity counts.
-Table 6: Results on ChineseLips, a real-world Chinese SlideASR dataset. The best and second-best results are in bold and underlined, respectively.
-[…]
-Contextless
-Qwen2-Audio 12.536
-Mi-Dasheng 3.311
-MiniCPM-o-2.6 2.252
-Qwen2.5-Omni-3B 1.937
-Qwen2.5-Omni-7B 2.243
-Qwen3-Omni-30B-A3B 2.202
-Slide text as context (Pipeline)
-Qwen2-Audio 84.291
-Mi-Dasheng 65.505
-Qwen3-Omni-30B-A3B 69.172
-Slide image as context (End-to-End)
-MiniCPM-o-2.6 76.203
-Qwen2.5-Omni-3B 24.847
-Qwen2.5-Omni-7B 14.340
-Qwen3-Omni-30B-A3B 41.930
-VAPO-3B (Ours) 1.548
-VAPO-7B (Ours) 1.298
+  • NE-FNR: The false negative ratio of named entities, calculated as $1-\frac{H}{N}$, where $H$ and $N$ denote the recognized and ground-truth entity counts.
 B.3 Baselines
@@
 The results show that our method achieves the lowest CER, demonstrating its effectiveness in real-world general-domain scenarios.
-Figure 5: Qualitative examples from SlideASR-R.
+Table 8: Results on ChineseLips, a real-world Chinese SlideASR dataset. The best and second-best results are in bold and underlined, respectively.
+[…]
+Contextless
+Qwen2-Audio 12.536
+Mi-Dasheng 3.311
+MiniCPM-o-2.6 2.252
+Qwen2.5-Omni-3B 1.937
+Qwen2.5-Omni-7B 2.243
+Qwen3-Omni-30B-A3B 2.202
+Slide text as context (Pipeline)
+Qwen2-Audio 84.291
+Mi-Dasheng 65.505
+Qwen3-Omni-30B-A3B 69.172
+Slide image as context (End-to-End)
+MiniCPM-o-2.6 76.203
+Qwen2.5-Omni-3B 24.847
+Qwen2.5-Omni-7B 14.340
+Qwen3-Omni-30B-A3B 41.930
+VAPO-3B (Ours) 1.548
+VAPO-7B (Ours) 1.298
 Appendix D Case Study
+Figure 7: Qualitative comparison on the SlideASR-R dataset.
+We compare the transcription results of the baseline (Qwen2.5-Omni-7B), a strong contextual baseline (Qwen3-Omni-30B-A3B with OCR), and our VAPO-7B.
+Red text indicates entity recognition errors (e.g., misrecognizing "CheXzero" as "checks zero"), while blue text highlights correctly transcribed entities.
 Figure 8: A failure case from SlideASR-R. OCR errors occurred due to low image resolution and small entity font size, leading to the loss of correct visual anchor points.
 D.1 Successful Case
-Table 7: Ablation results on the SlideSpeech dataset.
-ASR OCR VA Dev Test
-[…]
-12.22 12.74 5.26 95.17 19.99 20.71 9.80 94.44
-10.22 10.68 4.01 96.08 11.02 11.52 3.92 96.17
-9.97 10.49 3.83 96.30 11.74 12.25 4.47 95.63
-9.84 10.31 3.61 96.54 10.73 11.24 3.55 96.57
-Table 8: Ablation results on the SlideASR-S dataset.
-ASR OCR VA En Zh
-[…]
-100.08 53.19 18.72 86.86 65.62 9.62
-5.40 3.78 4.47 2.49 4.33 2.49
-4.98 3.71 4.23 2.58 4.54 2.82
-4.90 3.19 3.73 2.47 4.21 2.22
-Fig 5 shows a comparison of outputs from Qwen2.5-Omni-7B, and Qwen3-Omni-30B-A3B and our proposed VAPO-7B models on samples from the SlideASR-R dataset.
+Fig.
+7 shows a comparison of outputs from Qwen2.5-Omni-7B, and Qwen3-Omni-30B-A3B and our proposed VAPO-7B models on samples from the SlideASR-R dataset.
 Among them, Qwen2.5-Omni-7B uses audio-only input, Qwen3-Omni-30B-A3B uses OCR text extracted from the slide image as context, and VAPO-7B uses the slide image as context input.
@@
 D.2 Failure Case
-Fig 6 reveals a failure mode of our VAPO-7B model, originating from a visual perception error.
+Fig.
+8 reveals a failure mode of our VAPO-7B model, originating from a visual perception error.
 The issue begins in the <think> block, where the internal OCR component misidentifies a key entity.
@@
 However, this doesn’t diminish its overall advantage, as a contextless audio-only model would be equally, if not more, prone to failure when confronted with such inherent ambiguities in the source modalities.
-Table 9: Ablation results of reward function weights on SlideASR-Bench.
-Hyperparameter SlideASR-S (en) SlideASR-S (zh) SlideASR-R
-$\lambda_{1}:\lambda_{2}:\lambda_{3}:\lambda_{4}$ WER NE-WER NE-FNR WER NE-WER NE-FNR WER NE-FNR
-$1:1:1:1$ 4.90 3.19 3.73 2.47 4.21 2.22 27.28 19.31
-$1:1:1:2$ 5.27 3.34 3.78 2.50 4.30 2.09 27.67 17.73
-$1:1:2:1$ 5.32 4.12 3.91 2.48 4.38 2.09 30.35 22.17
-$1:2:1:1$ 5.17 3.45 3.80 2.51 4.23 1.99 27.54 21.18
+Table 9: Ablation results of functions on the SlideSpeech dataset.
+ASR OCR VA Dev set Test set
+[…]
+12.22 12.74 5.26 95.17 19.99 20.71 9.80 94.44
+10.22 10.68 4.01 96.08 11.02 11.52 3.92 96.17
+9.97 10.49 3.83 96.30 11.74 12.25 4.47 95.63
+9.84 10.31 3.61 96.54 10.73 11.24 3.55 96.57
+Table 10: Ablation results of the rewards on the SlideASR-S dataset.
+ASR OCR VA SlideASR-S (en) SlideASR-S (zh)
+[…]
+100.08 53.19 18.72 86.86 65.62 9.62
+5.40 3.78 4.47 2.49 4.33 2.49
+4.98 3.71 4.23 2.58 4.54 2.82
+4.90 3.19 3.73 2.47 4.21 2.22
+Appendix E More Ablation Results on Reward Functions
+Table 9 and Table 10 respectively present the ablation results of VAPO-3B (based on Qwen2.5-Omni-3B) on SlideSpeech Wang et al.
+(2024b) and SlideASR-S.
+The results indicate that different reward functions have a positive impact on the final performance.
+Appendix F More Cases of Attention Visualization
+Fig.
+9 further presents two cases of attention visualization.
+It can be seen that when transcribing key entities, the model is able to focus its attention on the same entities in the <think> block.
+This desirable property enables the model to accurately transcribe key entities in the speech.
 Table 11: Comparison of inference time and NE-FNR on the SlideASR-R dataset.
-Model Setting
-Inference time per sample (s)
-NE-FNR
+Model Setting Inference time per sample (s) NE-FNR
 Qwen3-Omni-30B-A3B Slide text as context 105.98 28.22
@@
 VAPO-7B Slide image as context 7.27 15.35
-Appendix E More Ablation Results
-E.1 Ablation Results of Reward Components
-Table 7 and Table 8 respectively present the ablation results of VAPO-3B (based on Qwen2.5-Omni-3B) on SlideSpeech Wang et al.
-(2024b) and SlideASR-S.
-The results indicate that different reward functions have a positive impact on the final performance.
-Figure 7: More cases of attention visualization.
-E.2 Sensitivity Analysis of Reward Weights
-To investigate the sensitivity of our VAPO method to the weights ($\lambda_{1}:\lambda_{2}:\lambda_{3}:\lambda_{4}$) of the four reward functions, we conducted an ablation study on the SlideASR-Bench.
-As shown in Table 9, we compared our default equal weighting scheme (1:1:1:1) against configurations where the weight of the OCR reward ($\lambda_{2}$), ASR reward ($\lambda_{3}$), or the Visual Anchoring reward ($\lambda_{4}$) was doubled.
-The results confirm that the 1:1:1:1 configuration provides the best overall performance, achieving the lowest Word Error Rate (WER) across all three subsets.
-This demonstrates the importance of a balanced approach to synergistic learning.
-The analysis also reveals a key trade-off.
-While doubling the Visual Anchoring reward (1:1:1:2) achieves the best NE-FNR (17.73) on the challenging real-world SlideASR-R set, it does so at the cost of higher overall WER.
-Conversely, overweighting the ASR reward (1:1:2:1) proved counterproductive, as it discourages the model from leveraging crucial visual context and significantly degrades performance.
-Therefore, we conclude that an equal weighting of the reward functions is a simple and effective choice for the VAPO method.
-This configuration was used for all other experiments reported in this paper.
-Appendix F More Cases of Attention Visualization
-Figure 6 further presents two cases of attention visualization.
-It can be seen that when transcribing key entities, the model is able to focus its attention on the same entities in the <think> block.
-This desirable property enables the model to accurately transcribe key entities in the speech.
 Appendix G Inference Latency Analysis
@@
 While not yet suitable for real-time applications, the latency of VAPO is a reasonable trade-off for its state-of-the-art performance, particularly for offline transcription tasks where accuracy is paramount.
+Figure 9: More cases of attention visualization.